### v5

# Liu, Rahimian, and Yu

[…]
Liu, Rahimian, and Yu
[…]
Seeding with Differentially Private Network Information
[…]
Seeding with Differentially Private Network Information
[…]
In public health interventions such as distributing preexposure prophylaxis (PrEP) for HIV prevention, decision makers often use seeding algorithms to identify key individuals who can amplify intervention impact. However, building a complete sexual activity network is typically infeasible due to privacy concerns. Instead, contact tracing can provide influence samples—observed sequences of sexual contacts—without full network reconstruction. This raises two challenges: protecting individual privacy in these samples and adapting seeding algorithms to incomplete data. We study differential privacy guarantees for influence maximization when the input consists of randomly collected cascades. Building on recent advances in costly seeding, we propose privacy-preserving algorithms that introduce randomization in data or outputs and bound the privacy loss of each node. Theoretical analysis and simulations on synthetic and real-world sexual contact data show that performance degrades gracefully as privacy budgets tighten, with central privacy regimes achieving better trade-offs than local ones.
[…]
social networks, differential privacy, influence maximization, sample complexity, approximation algorithms, sexual behavior, HIV, public health
1 Introduction
When designing interventions in public health, development and education, decision makers rely on data from social networks to target a small number of people, leveraging peer effects and social contagion to deliver the greatest welfare benefits to the population. Many of these seeding methods are implemented without complete knowledge of social network structures—for example, by targeting nominated friends of random individuals (kim2015social) or random pairs of friends (alexander2022algorithms)—and enhance adoption of public health interventions by driving indirect treatment effects among people who were affected by spillovers from others’ treatment (even though they were not part of the initial treatment group). However, even without full knowledge of the underlying network, individuals who participate in partial social network data collection—especially members of minority or stigmatized groups—still face privacy risks. For instance, in the context of HIV intervention, if contact tracing begins with individuals who have already tested positive, others may be reluctant to participate in the survey due to the perceived risk of being associated with high-risk contacts—such as being mistakenly identified as part of a transmission chain, facing social stigma, or having sensitive health information inferred from their network connections. At the same time, social network information plays a critical role in designing effective HIV prevention strategies (Wilder_Onasch-Vera_Diguiseppi_Petering_Hill_Yadav_Rice_Tambe_2021). For example, the distribution of pre-exposure prophylaxis (PrEP) can benefit from seeding algorithms that identify high-risk individuals based on data from social networks (ross2016implementation). However, constructing a complete sexual activity network for this purpose is challenging due to the sensitivity of the sexual behavior data. Instead, contact tracing can provide influence samples, which capture sequences of sexual activity without requiring full network knowledge. This introduces two key challenges: protecting the privacy of individuals contributing to contract tracing data (i.e., influence samples) and adapting seeding algorithms to effectively leverage these partial data. Developing privacy-preserving methods for network data collection and targeted interventions is critical for responsible public health interventions in social networks, particularly in vulnerable communities such as homeless youth, who are estimated to have a six to twelve times higher risk of HIV than housed youth (caccamo2017narrative). Such methods ensure that privacy concerns do not undermine the trust and community engagement necessary for successful public health initiatives.
In this paper, we propose randomization techniques based on the differential privacy formalism to provide social network participants with plausible deniability. Under these schemes, network participants can make a credible claim that their associates are targeted (e.g., recruited into treatment) by random choice, irrespective of their associations. We investigate how the targeting efficacy is affected by decreasing the privacy budget that entails more noise in the collected data (with local privacy mechanisms) or the targeting algorithm’s output (with central privacy mechanisms).
After summarizing our main contributions in Section 1.1 and providing detailed comparisons with related works in Section 1.2, in Section 2 we provide preliminaries including hardness results that demonstrate the inadequacy of mainstream approaches to the privacy of social network data for seeding interventions and motivate our proposed notion of “influence sample differential privacy”. In Section 3, we provide privacy mechanisms for seeding in central (Section 3.1) and local (Section 3.2) regimes. In Section 4, we test the performance of our privacy mechanisms on empirically-grounded networks of men who have sex with men (MSM). We end with a discussion of limitations, future directions, and the concluding remarks in Section 5. Complete proofs and supplementary results are provided in several appendices at the end.
1.1 Summary of the main results
Differential privacy (DP) offers an information-theoretic foundation for privacy by focusing on protecting input data against revelations of algorithmic outputs (dwork2014algorithmic). In mathematical formulation, a randomized mechanism $\mathcal{M}$ is $\epsilon$-DP if for any two adjacent input datasets $D$ and $D^{\prime}$ in the input domain and any subset $A$ in the output range we have:
$\frac{\Pr[\mathcal{M}(D)\in A]}{\Pr[\mathcal{M}(D^{\prime})\in A]}\leq e^{\epsilon}.$ (1)
The $\epsilon$-DP criterion in Equation 1 limits the inferences of any observer/adversary about the input data based on the algorithm output. The notion of adjacency between datasets is typically defined so that neighboring datasets differ in contributions of an individual entity, and thus $\epsilon$-DP protects individuals’ contributions from ramification of an adversarial inference. Hence, different notions of adjacency will imply different protections for individuals. In the sequel, we explain our proposed notion of adjacency for seeding with randomized contagion data (which we term “influence samples”) and the nature of privacy protections that it offers.
Seeding with partial and private network information.
We study a classical, NP-hard network optimization problem, known as “influence maximization”, which asks to select $k$ nodes on a graph of size $n$ such that the expected spread size from seeding those $k$ nodes is maximum in a randomized network diffusion model, whereby each edge passes on the influence with an independent cascade probability $p$. Given the social network graph $G$ and an integer $k\geq 1$, we denote the expected spread size from infecting the nodes in $S$ by $I_{G}(S)$ and write $\mbox{\sf OPT}=\max_{S:|S|\leq k}I_{G}(S)$ for the optimum solution to the $k$-influence maximization problem. It is an example of a cardinality-constrained combinatorial optimization on a graph with a monotone, submodular objective that admits a tight (11/e)(1-1/e) approximation guarantee using the greedy node selection algorithm. This problem has occupied a central place in the graph and social networks algorithms literature since its introduction by kempe2003maximizing. There is a wealth of literature (spanning almost two decades) on efficient approximations of this problem (kempe2003maximizing, Kempe2015-pi, chen2009efficient, borgs2014maximizing, du2024adaptive). However, the graph structure of the input data has stunned the design of DP influence maximization algorithms. Common operationalizations of DP for graph algorithms, such as node or edge DP, define adjacent inputs in Equation 1 based on neighboring graphs differing in a node or an edge. However, this adjacency criterion is often too stringent for network-based problems whose outputs are highly sensitive to such structural changes. Even the removal or addition of one node or edge can substantially alter the entire contagion process, leading to drastic changes in the resulting diffusion outcomes (see Section 1.2 and Section 7.3 for detailed citations and comparisons). Our main hardness results indicate that it is impossible to obtain useful performance guarantees for influence maximization subject to node- or edge-DP constraints (see Section 2.2, Propositions 2.5 and 2.6).
A new privacy notion for social network data collection and intervention design.
Our main contribution is to formulate a new differential privacy notion—influence-sample differential privacy (ISDP)—that provides guarantees for influence maximization when the social network is unknown and the algorithm operates on randomly collected samples of past diffusion cascades, or influence samples, rather than the entire network graph. Building on prior work on seeding with costly network information (costlyseeding) and the sample complexity of influence maximization (sadeh2020sample, borgs2014maximizing), our framework protects the privacy of individuals whose behavioral data are used for intervention design.
Influence Sample Differential Privacy (ISDP).
We define $\epsilon$ influence sample differential privacy ($\epsilon$-ISDP)” by controlling the odds ratio of the output of the seeding algorithm on adjacent influence samples, i.e., those differing in participation of a single node in a single influence cascade.
Influence samples operate at the activity level, capturing who participated in which diffusion events and thereby potentially revealing whether a person had direct or indirect contact with high-risk individuals. This naturally raises privacy concerns of using influence samples. The activity-level notion of privacy can be understood as a form of record-level or action-level differential privacy, in which each participation record—representing an individual’s involvement in a diffusion event—is treated as an independent privacy unit (liu2024cross, messing2state). Under $\epsilon$-influence-sample differential privacy ($\epsilon$-ISDP), two datasets are considered adjacent if they differ in the participation of a single individual in a single cascade—equivalently, if an individual’s participation in a specific event (e.g., attending a party, a meeting, or visiting a physical location identified during contact tracing) changes. The privacy parameter $\epsilon$ bounds the extent to which the algorithm’s output can vary between adjacent datasets, ensuring that the released seed set discloses only limited information about any individual’s participation in different activities. Smaller values of $\epsilon$ correspond to stronger privacy guarantees, offering greater protection against inferences about private behaviors.
To illustrate the type of information that ISDP protects, consider a contact-tracing scenario during a disease outbreak such as COVID-19. Investigators may identify several social gatherings or public events—such as parties, concerts, or community meetings—as potential starting points of infection. Each tracing record then corresponds to one such event and indicates who attended it, forming a partial view of how the infection might have spread through direct or indirect contacts. While these data are essential for designing effective interventions, they are also highly sensitive: participation in a single event record could reveal a person’s association with potentially infected individuals or their other political, personal or professional orientations that are associated with the event. An adversary observing the final intervention targets could thus infer whether someone appeared in these event records. ISDP mitigates this risk by bounding the influence of any single event record on the seeding outcome, ensuring that adding or removing one person’s participation in an event does not substantially alter the recommended targets. In doing so, ISDP gives individuals the ability to plausibly deny their attendance at high-risk events, while still enabling accurate population-level intervention design.
Sample complexity with differential privacy and approximation guarantees.
We analyze ISDP under two standard privacy regimes—central and local differential privacy—which differ in the level of trust placed in the curator. Under central DP, a trusted curator accesses the raw cascade data and enforces privacy by perturbing the algorithm’s output. Under local DP, individuals randomize their own influence samples before submission, thereby protecting privacy at the data-collection stage.
To achieve these guarantees, we design two efficient $\epsilon$-ISDP seeding algorithms based on the exponential mechanism (mcsherry2007mechanism) and the randomized response mechanism (warner1965randomized), corresponding to the central and local privacy regimes, respectively. We derive performance bounds in terms of sample size $m$, network size $n$, seed set size $k$, and privacy budget $\epsilon$. Our results show that the central $\epsilon$-ISDP algorithm achieves near-optimal influence, infecting at least $(1-1/e)\mbox{\sf OPT}-\alpha n$ nodes with high probability, and requires about $m=\max\{\tfrac{12}{\alpha\epsilon},\tfrac{9}{\alpha^{2}}\}k\ln n$ influence samples. In contrast, the local $\epsilon$-ISDP algorithm needs substantially more samples due to the additional noise introduced by local randomization.(see details in Section 3, LABEL:thm:exp:acc and LABEL:thm:rr:acc).
1.2 Comparisons with related work
Seeding network diffusions.
A long line of literature, going back to the seminal works of kempe2003maximizing, Kempe2015-pi, proposes efficient approximation algorithms for influence maximization based on the submodularity of the influence function (chen2009efficient, wang2012scalable).This problem has been widely applied in marketing and public health to promote the adoption of new products and interventions through social networks (han2023cost, kahr2024impact). For example, banerjee2013diffusion demonstrate that network-based seeding substantially improves participation in social programs, while kim2015social show that targeting key individuals can increase community adoption of water purification methods in field experiments. In the context of public health, social network interventions have been used to increase awareness of HIV prevention among homeless youth (yadav2018bridging) and to reduce bullying among adolescents (paluck2016changing). In all such settings, collecting complete social network data is difficult, costly, or infeasible, and poses critical privacy risks to individuals. In this paper, we focus on providing differential privacy guarantees for influence maximization when the social network is unknown and influential nodes must be inferred from limited and potentially sensitive diffusion samples.
Social network data privacy.
Much of the previous work on social network data privacy focuses on the identifiability of nodes from released graph data, with or without additional information on node attributes and their neighborhoods (romanini2021privacy). Classical anonymization techniques mask the attributes of the nodes and perturb, modify, randomize, or aggregate the structure of the graph to prevent reidentification of the nodes within a confidence level, typically achieving $k$ anonymity (wu2010survey, zheleva2011privacy). However, statistical disclosure control of collected data requires safeguarding data donor’s privacy against adversarial attacks where common anonymization techniques are shown to be vulnerable to various kinds of identification (barbaro2006face), linkage and cross-referencing (narayanan2008robust, sweeney2015only), statistical difference and reidentification attacks (kumar2007anonymizing). However, while edge sampling at data collection can somewhat mitigate the risk of reidentification of nodes (romanini2021privacy, jiang2021applications), seeded nodes can reveal additional side information about who has contributed to data collection. For example, once a community leader is chosen, his or her close contacts may suffer privacy leaks even though no social network data is published. Our privacy guarantees for influence maximization follow the DP formalism to provide network participants and data donors with plausible deniability with regard to their contribution to the collected data or their role in shaping the algorithm output (the seeded nodes). DP is a gold standard for privacy-preserving data analysis and avoids the above drawbacks by abstracting away from the attack models and controlling the privacy leak at an information-theoretic level according to Equation 1.
DP analysis of graphs.
Differential privacy (DP) provides a general framework for designing privacy-preserving algorithms and has been adopted in large-scale applications such as the US Census (garfinkel2022differential) and major technology platforms (erlingsson2014learning, appleDP). It has been successfully applied to a wide range of optimization problems, including combinatorial (gupta2010differentially), submodular (pmlr-v70-mitrovic17a), and online submodular optimization (salazar2021differentially). Existing work on differential privacy for graphs—such as edge DP (nissim2007smooth, eden2023triangle) and node DP (kasiviswanathan2013analyzing, raskhodnikova2016lipschitz, kalemaj2023node)—has primarily focused on protecting global graph statistics or subgraph counts (blocki2022privately, hay2009accurate). However, these approaches are not directly applicable to influence maximization, and our hardness results show that node or edge DP is often too stringent to yield meaningful performance guarantees. Similar limitations have been observed for other graph algorithms that are highly sensitive to the addition or removal of individual nodes or edges (gupta2010differentially, epasto2022differentially). To address these limitations, we introduce influence sample differential privacy (ISDP), which shifts the input from an explicit social network to a collection of randomly sampled diffusion cascades. ISDP bounds information leakage about each node’s participation in these cascades—either at the data level (local ISDP) or through the algorithm’s output (central ISDP). While local ISDP can be achieved by randomizing individual participation reports, coordinating such randomization in a decentralized and context-aware manner remains challenging (imola2021locally, dhulipala2022differential).
2 Problem Setup and Preliminary Results
2.1 Influence sampling and maximization
In the independent cascade (IC) model, infections spread via an edge-weighted undirected graph $G=(V,p)$ on $n$ nodes. The set of nodes $V$ is indexed by $n]:={1,\ldots,n}$, and $p:V\times V\to[0,1]$ assigns probability weights to every edge in $G$. The infection starts from a seed set $S$ and propagates through the network in discrete rounds. Each newly infected node has a single opportunity to activate each of its neighbors. For every undirected edge $e=(v,u)$, the transmission probability $p_{e}=p(v,u)=p(u,v)\in[0,1]$ represents the chance that the infection spreads from $v$ to $u$ (or $u$ to $v$) in the next round after one endpoint becomes infected.
Following kempe2003maximizing, this stochastic diffusion process can be viewed equivalently as sampling a random realization of the network. Specifically, we draw an unweighted graph $g$ from $G$ by including each edge $e$ independently with probability $p_{e}$. The eventual set of infected nodes is precisely the set of nodes reachable from the seed set $S$ in $g$. This reachability-based formulation offers a convenient way to analyze influence propagation and will be used throughout the paper.
Influence function.
For a given realization $g\sim{\mathcal{G}}$ and a seed set $S\subseteq V$, let $C_{g}(S)$ denote the set of nodes reachable from $S$ in $g$, where $\mathcal{G}$} denotes the distribution of subgraphs randomly drawn under the IC model. The influence of $S$ on $g$ is defined as $I_{g}(S)=|C_{g}(S)|$, the number of nodes activated by $S$. Its expectation over random graph realizations, $I_{G}(S):=\operatorname*{\mathbb{E}}_{g\sim{\mathcal{G}}}[I_{g}(S)]$, measures the expected spread of $S$ under the IC model. We also define the conditional reachable set $C_{g}(S|T):=C_{g}(S)\setminus C_{g}(T)$, representing nodes reachable from $S$ but not from $T$, and its expected size $I_{G}(S|T):=\operatorname*{\mathbb{E}}_{g\sim{\mathcal{G}}}[|C_{g}(S|T)|]$.
The influence maximization problem.
Given graph $G$ and integer $k\geq 1$, the $k$-influence maximization problem is to find a set $S\subseteq V$ of at most $k$ nodes maximizing the expected number of reachable nodes, $I_{G}(S)$. We write $\mbox{\sf OPT}=\max_{S:|S|\leq k}I_{G}(S)$ for the maximum expected influence of any set of $k$ nodes.
Influence samples under limited information.
While there exists extensive literature on developing fast algorithms for maximizing influence when the contagion model $G=(V,p)$ is known, in many real-world settings the underlying network or edge probabilities are unobservable. Learning $G$ directly may be too costly or even infeasible due to privacy restrictions or incomplete data. To address this limitation, we consider a query model where the algorithm observes only influence samples—random realizations of who could potentially reach whom in the diffusion process.
Intuitively, each influence sample corresponds to one instance of data collection that records individuals’ participation in a particular activity—such as attending a party or visiting a specific location. Conducting such surveys across multiple, independently selected activities yields multiple influence samples. Each sample thus represents an independent observation of how individuals participate in diffusion-related events.
Formally, we sample a target node $u\in V$ uniformly at random and a graph realization $g\sim{\mathcal{G}}$, and define the influence sample as a binary vector $x=(x_{1},\ldots,x_{n})^{\top}\in\{0,1\}^{n}$ where $x_{v}=\mathbf{1}[\,u\in C_{g}(v)\,]$, $x_{v}=1$ if node $v$ can reach (or influence) the target node $u$ in $g$. We can equivalently view $x$ as the subset of nodes $x=\{v\in V:x_{v}=1\}$. A collection of $m$ independent samples is denoted by ${\mathbf{x}}=(x^{1},\dots,x^{m})\in\{0,1\}^{n\times m}$, where each $x^{t}$ corresponds to one realization.
For simplicity, we identify each influence sample $x$ with its corresponding set of active nodes and use standard set notation. For any subset $S\subseteq V$, $S\cap x$ denotes the set of nodes in $S$ that appear in the sample $x$. When referring to a single node $v\in V$, we write $S\cap v$ as shorthand for $S\cap\{v\}$. This convention will be used throughout the paper.
Influence sample greedy seeding algorithm.
How can we design a seeding algorithm $M$ that takes $m$ influence samples $\mathbf{x}$} and produces a set of $k$ nodes? Here we present a seeding algorithm by costlyseeding. The core idea of Algorithm 1 is to estimate the influence function from the influence samples and to run the greedy algorithm. Specifically, given $m$ realized influence samples $\mathbf{x}$} and a subset of nodes $S$ in $G$, we define the $S$-reachable influence samples as: $C_{\mathbf{x}}(S):=\{t\in[m]:\exists v\in S,x_{v}^{t}=1\},$ which is the set of influence samples whose activated nodes are reachable from the set $S$, and $C_{\mathbf{x}}(S|T)=C_{\mathbf{x}}(S)\setminus C_{\mathbf{x}}(T)$. Alternatively, if we treat $\mathbf{x}$} as a collection of $m$ subsets $x^{1},\dots,x^{m}$, $C_{\mathbf{x}}(S)=\{t:S\cap x^{t}\neq\emptyset\}$. Finally, we define $I_{\mathbf{x}}(S)=\frac{n}{m}|C_{\mathbf{x}}(S)|$ and $I_{\mathbf{x}}(S|T)=\frac{n}{m}|C_{\mathbf{x}}(S|T)|$. The following proposition states that the expectation of $I_{\mathbf{x}}$ is equal to $I_{G}$. We will use $v$ instead of $\{v\}$ for singleton sets to simplify the notation, for example, writing $I_{G}(v)$ for $I_{G}(\{v\})$.
Proposition 2.1 (Observation 3.2. of DBLP:journals/corr/abs-1212-0884 – cf. borgs2014maximizing)
Given a positive integer $m$, a graph $G=(V,p)$, and sets of nodes $S,T\subseteq V$, if $\mathbf{x}$} is the random variable that encodes $m$ influence samples, then $\operatorname*{\mathbb{E}}_{{\mathbf{x}}}[I_{{\mathbf{x}}}(S)]=I_{G}(S)$, and $\operatorname*{\mathbb{E}}_{\mathbf{x}}[I_{\mathbf{x}}(S|T)]=I_{G}(S|T)$. Additionally, by Hoeffding’s inequality, for all $\epsilon>0$ we have: $\Pr[|I_{{\mathbf{x}}}(S)-I_{G}(S)|\geq\epsilon n]\leq 2\exp(-2\epsilon^{2}m),$ where the probability is over the randomness in the generative model of influence samples $x$.
With the above observation, costlyseeding design a seeding algorithm that only uses influence samples and show that if the number of influence samples is of the order $k^{2}(\log n/\alpha)/\alpha^{3}$, then Algorithm 1 is optimal up to the additive error $\alpha n$. Informally, Algorithm 1 is a greedy algorithm that sequentially queries Algorithm 2 for the seed that maximizes marginal improvement. In this paper, we will follow a similar idea that combines a greedy algorithm and an estimation algorithm for marginal improvement.
Algorithm 1 Influence sample greedy seeding algorithm
1:$m$ influence samples $\mathbf{x}$} and total number of seeds $k$.
2:Set $S_{0}=\emptyset$
3:for $i$ from $1$ to $k$ do
4:  $v_{i}=\mathcal{M}({\mathbf{x}},S_{i-1})$ by Algorithm 2
5:  $S_{i}=S_{i-1}\cup v_{i}$
6:end for
7:Return $S_{k}$
Algorithm 2 Single seed selection, $\mathcal{M}({\mathbf{x}},S)$
1:$m$ influence samples $\mathbf{x}$} and a seed set $S$.
2:for $v\in[n]\setminus S$ do $\triangleright$ Chooses the $|S|+1$-th seed.
3:  Compute $I_{\mathbf{x}}(v|S)=\frac{n}{m}|C_{\mathbf{x}}(S\cup v)\setminus C_{\mathbf{x}}(S)|$.
4:end for
5:Return $argmax_{v\in[n]\setminus S}I_{\mathbf{x}}(v|S)$
In Algorithm 1, we modify the baseline influence-sample-based greedy algorithm to sequentially call Algorithm 2 for selecting each seed. Instead of collecting fresh influence samples at every greedy step, Algorithm 1 reuses the existing influence samples, thereby achieving substantially better sample efficiency. We formally establish this improvement in LABEL:thm:exp:acc.
Theorem 2.2 (label = thm:original, name = Theorem 2 of costlyseeding)
Given the graph $G$ on $n\geq 2$ nodes, we want to seed $k\leq n$ nodes. For any $0<\alpha\leq 1$, the output of Algorithm 1 can infect $(1-1/e)\mbox{\sf OPT}-\alpha n$ nodes using no more than $81k^{2}\log(6nk/\alpha)/\alpha^{3}$ influence samples.
It is well known that even with complete knowledge of the graph $G$, no efficient algorithm can perform better than $(1-1/e)\mbox{\sf OPT}$ unless P equals NP. costlyseeding show that without the knowledge of $G$, Algorithm 1 can approach an optimal efficient algorithm when the number of influence samples is large enough. In this paper, we ask if we can achieve similar performance when privacy constraints are imposed. costlyseeding also show that one cannot hope to provide a constant factor approximation guarantee using $o(n)$ influence samples.
Theorem 2.3 (Theorem 3 of costlyseeding)
Let $0<\alpha<1$ be any constant. For any algorithm with $o(n)$ samples, there exists a weighted graph $G$ so that the output of the algorithm can only infect fewer than $\alpha\mbox{\sf OPT}$ nodes.
2.2 Influence Sample Differential Privacy
Intuitively, a differentially private (DP) algorithm ensures that the algorithms output on two adjacent inputs are close to each other, making them difficult to distinguish. However, there are multiple alternative definitions of the adjacency relation in graph-structured input data. One classical notion is node DP, where two graphs are adjacent if they are different by the connections of one node. Another is edge DP, where two graphs are adjacent if they differ by one edge (hay2009accurate). While these definitions are well established, they often require adding excessive noise to ensure privacy. Moreover, they are ill-suited for sensitive domains such as public health contact tracing, where the underlying network is unclear and constructing an explicit network may be infeasible. In practice, contact tracing relies on surveys from confirmed cases rather than on full network data, both for practical and privacy reasons. Given the contextual and dynamic nature of contact data, a fixed network representation is often poorly defined. We therefore abstract away from the network and work directly with influence samples, which capture how contact patterns propagate over time without requiring complete network information. A more detailed discussion of the relationship between node- or edge-level DP and ISDP is provided in Section 7.3.
We study algorithms that only use these influence samples, where two collections of $m$ influence samples, $\mathbf{x}$} and $\mathbf{x}$$^{\prime}$ in $\{0,1\}^{n\times m}$, are “influence sample” adjacent if one entry can be modified to transform $\mathbf{x}$} into $\mathbf{x}$$^{\prime}$. We denote this adjacency by ${\mathbf{x}}\underset{is}{\sim}{\mathbf{x}}^{\prime}$. Given this adjacency relation on influence samples, we define a stochastic seeding algorithm $M$ to be DP if the algorithm’s output does not change significantly when applied to adjacent influence samples. In the context of HIV prevention, the seed set output by the algorithm—corresponding to individuals recommended for pre-exposure prophylaxis (PrEP)—is typically observable in practice, because healthcare workers must implement the intervention for those selected individuals. doi:10.1073/pnas.1510612113 note that when performing social network search algorithms, it is important to distinguish between entities that require privacy protection and those that do not. The seed set, which represents the actionable output of the algorithm, falls into the latter category: it must be disclosed to enable implementation and thus cannot itself be protected under differential privacy. Instead, the privacy guarantee should apply to the process of identifying the seed set—specifically, to individuals’ participation in diffusion samples and their network relationships. Consequently, an adversary may still attempt to infer, from differences between adjacent influence samples, whether a specific individual is reachable from a high-risk participant, potentially revealing sensitive contact information and exposing them to social stigma or discrimination.
Definition 2.4 (ISDP)
Given $\epsilon\geq 0$, $n,m\geq 1$, and $k\in[n]$, a function $M^{k}:\{0,1\}^{n\times m}\to\mathcal{Y}$ is $\epsilon$-influence sample differentially private ($\epsilon$-ISDP, in short) if for all outputs ${\mathbf{y}}\in\mathcal{Y}$ of $k$ seeds, and all pairs of adjacent datasets (collection of influence samples), $\mathbf{x}\underset{is}{\sim}\mathbf{x}^{\prime}$, we have: $\Pr[M^{k}(\mathbf{x})={\mathbf{y}}]\leq e^{\epsilon}\Pr[M^{k}(\mathbf{x}^{\prime})={\mathbf{y}}].$
We believe that the study of influence sample DP is more natural and practical than conventional node and edge DP for seeding algorithms (kasiviswanathan2013analyzing, raskhodnikova2016lipschitz). First, the notion of node differential privacy (DP) allows an agent to misreport or withhold her social connections. However, in the context of network cascade data, an individual has limited ability to alter her network ties to affect the contagion dynamics; instead, she can decide whether to acknowledge or deny participation in a particular activity, such as denying that she visited a specific place in a contact-tracing survey. Therefore, we argue that defining adjacent datasets based on influence samples as input, rather than raw social network data, is more natural and realistic for contact tracing and public health applications. Moreover, we show in Propositions 2.5 and 2.6 that node DP and edge DP are too strict for influence maximization and hardly admit any utility guarantees.
Given two edge-weighted undirected graphs $G=(V,p)$ and $G^{\prime}=(V^{\prime},p^{\prime})$, we say $G$ and $G^{\prime}$ are node adjacent if $V=V^{\prime}$ and there exist $v\in V$, such that $p_{e}=p_{e}^{\prime}$ for all $e\notniFromTxfonts v$. Similarly, $G$ and $G^{\prime}$ are edge adjacent if $V=V^{\prime}$ and there exists $u,v\in V$ such that $p_{e}=p_{e}^{\prime}$ for all $e\neq(u,v)$. For $k<n$, we say that a seeding algorithm $M^{k}$ is $\epsilon$-node (or edge) DP if for any $k$, node (or edge) adjacent graphs $G,G^{\prime}$, and output $\mathbf{y}$}, $\Pr[M^{k}(G)={\mathbf{y}}]\leq e^{\epsilon}\Pr[M^{k}(G^{\prime})={\mathbf{y}}]$.
The following proposition (proved in Section 6.1) shows that seeding algorithms with node DP guarantees cannot produce high-quality seeds.
Proposition 2.5 (Hardness of Influence Maximization under Node DP)
For any $\epsilon$-node DP seeding algorithm $M^{k}(G)$ with $n>1$ and $\epsilon>0$, there exist graphs $G$ on $n$ nodes for which the following holds. When the maximum degree of $G$ is n1n-1, $\max_{S:|S|=1}I_{G}(S)-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\geq\frac{n}{4}-(e^{\epsilon}+1)$. When the maximum degree satisfies $2\leq d<n-1$, $(1-1/e)\max_{S:|S|=1}I_{G}(S)-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\geq\frac{1}{10e^{\epsilon}}\Bigl(1-\frac{d}{d-1}\Bigl(\frac{1}{e}+\frac{1}{10}\Bigr)\Bigr)n-1+\frac{d}{d-1}\Bigl(\frac{1}{e}+\frac{1}{10}\Bigr)$.
When the maximum degree of $G$ is n1n-1, we use the star graph, since it differs from a totally isolated graph by the connection of a single node; a node-DP algorithm must output similar seed sets on these two graphs and thus yields low-quality seeds. When the maximum degree satisfies $2\leq d<n-1$, we construct $G^{\prime}=(V,p^{\prime})$ on $n$ nodes consisting of $l$ disjoint d1d-1-regular trees $S_{1},\dots,S_{l}$, each with roughly $n/l$ nodes and edges $p^{\prime}_{e}=1$ within trees and $p^{\prime}_{e}=0$ otherwise, and obtain $G=(V,p)$ by adding one node $v$ that connects the root nodes of two trees $S_{i}$ and $S_{j}$. Under node DP, the algorithm must produce outputs with similar probabilities on $G$ and $G^{\prime}$, so it cannot assign substantially higher probability to the high-quality seeds in $G$ than in $G^{\prime}$, where such seeds are ineffective, leading to low expected influence. Based on these constructions, the following proposition (proved in Section 6.2) further shows that no edge-DP seeding algorithm can output high-quality seeds either, and both hardness results extend to the approximate $(\epsilon,\delta)$-DP setting where Equation 1 allows an additive $\delta$; see Section 6.2.
Proposition 2.6 (Hardness of Influence Maximization under Edge DP)
For any $\epsilon$-edge DP seeding algorithm $M^{k}(G)$ with $n>1$ and $\epsilon>0$, there exists a graph $G$ on $n$ nodes whose maximum degree satisfies $d\geq 2$ such that $(1-1/e)\max_{S:|S|=1}I_{G}(S)-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\geq\frac{1}{10e^{\epsilon}}\Bigl(\frac{4}{5}-\frac{2}{e}\Bigr)n-\frac{4}{5}+\frac{2}{e}.$.
Compared to LABEL:thm:original, the above results show that node-DP (or edge-DP) algorithms cannot perform close to $(1-1/e)\mbox{\sf OPT}$ even with full knowledge of $G$. In Section 6.2 (Proposition 6.1), we show that similar lower bounds hold even if we consider additive relaxations of the node and edge DP, allowing the $\Pr[M^{k}(G)={\mathbf{y}}]\leq e^{\epsilon}\Pr[M^{k}(G^{\prime})={\mathbf{y}}]$ condition to be violated up to a $\delta$ additive factor. Building on the above hardness results for node- and edge-level DP, we next design ISDP algorithms that achieve performance comparable to Algorithm 1.
3 Differentially Private Seeding Algorithms
In this section, we provide two DP seeding algorithms that satisfy the DP requirements either centrally or locally. First in Section 3.1, we use an exponential mechanism that ensures that the seed selection output is centrally DP. Next in Section 3.2, we use the randomized response algorithm to achieve local DP according to the criteria set forth in Definition 2.4. LABEL:thm:exp:acc and LABEL:thm:rr:acc provide our sample complexity bounds in each case to achieve $\epsilon$-ISDP with a $(1-1/e)\mbox{\sf OPT}-\alpha n$ performance guarantee. In particular, we observe that the sample complexity bound for achieving local DP is much harsher at $O(k^{3}\epsilon^{-2k^{2}}\ln n/\alpha^{2})$ compared to $O(k\ln n/(\alpha\epsilon))$ for central DP. This exponential separation of the dependency of sample complexity on $k$ is reflected in our simulation results in Section 4, where the exponential mechanism shows a faster improvement in performance with an increasing number of influence samples and its output has a significantly higher expected spread size when the privacy budget ($\epsilon$) is low.
3.1 Central DP with Exponential Mechanism
In this section, we design a differentially private (DP) seeding algorithm based on the exponential mechanism, building upon and modifying Algorithm 1. To ensure privacy protection, we replace the seed selection step in Algorithm 2 with an exponential mechanism that selects a seed node according to the softmax distribution induced by the scores under differential privacy. Formally, to achieve $\epsilon$-DP, let the output set be $\mathcal{V}$ and the scoring function be $u:\{0,1\}^{n\times m}\times\mathcal{V}\to\mathbb{R}$ with sensitivity $\Delta(u):=\max_{v\in\mathcal{V},\,{\mathbf{x}}{\sim}{\mathbf{x}}^{\prime}}|u({\mathbf{x}},v)-u({\mathbf{x}}^{\prime},v)|$; the exponential mechanism $\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}},S)$ selects $v\in\mathcal{V}$ with probability proportional to $\exp\!\big(\epsilon\,u({\mathbf{x}},v)/(2\Delta(u))\big)$ (dwork2014algorithmic). In our setting, $\mathcal{V}=[n]\setminus S$ and $u({\mathbf{x}},v)=I_{\mathbf{x}}(v\mid S)=\tfrac{n}{m}\,\big|C_{\mathbf{x}}(S\cup\{v\})\setminus C_{\mathbf{x}}(S)\big|$, where $C_{\mathbf{x}}(S)$ denotes the set of influence samples whose activated nodes are reachable from the seed set $S$; the sensitivity is $\Delta(u)=\tfrac{n}{m}$. Given $\epsilon>0$ and an integer $k$, we define the DP seeding algorithm $\mathcal{M}^{k,\epsilon}_{\mathrm{exp}}({\mathbf{x}})$, which runs Algorithm 1 but replaces each single-seed selection step of Algorithm 2 with its DP variant $\mathcal{M}^{\epsilon}_{\mathrm{exp}}({\mathbf{x}},S)$ in Algorithm 3, i.e., an exponential-mechanism selection using $I_{\mathbf{x}}(\cdot\mid S)$ as the score.
Algorithm 3 $\epsilon$-ISDP exponential mechanism, $\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}},S)$
1:$m$ influence samples ${\mathbf{x}}=(x^{1},\dots,x^{m})^{\top}$, a seed set $S$, and privacy budget $\epsilon\geq 0$.
2:for $v\in[n]\setminus S$ do
3:  Compute $I_{\mathbf{x}}(v|S)$
4:end for
5:Return $v$ with probability proportional to $\exp\left(\frac{\epsilon m}{2n}I_{\mathbf{x}}(v|S)\right)$
Note that a naive algorithm that runs the exponential mechanism directly on all $k$ seeds at once is not efficient because the size of the output space is combinatorially large $\Omega(n^{k})$. To ensure that our algorithm is efficient, we sequentially run the exponential mechanism $k$ times to estimate and choose seeds similar to the greedy algorithm of costlyseeding, but reducing their time complexity to $O(nmk)$. LABEL:thm:exp:inf shows that our modified greedy algorithm ($\mathcal{M}_{exp}^{k,\epsilon}({\mathbf{x}})$) that applies the exponential mechanism ($\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}},S)$) for single seed selection $k$ times, is $\epsilon$-ISDP following Definition 2.4, and also provides a tail bound on the approximation error for maximization of $I_{\mathbf{x}}$ using $\mathcal{M}_{exp}^{k,\epsilon}({\mathbf{x}})$. Its proof in Section 6.3 follows from the facts that $I_{\mathbf{x}}$ is a submodular function, and the exponential mechanism in Algorithm 3 can be seen as a noisy maximum function (softmax) that implements an approximate greedy step in LABEL:thm:exp:utility.
Theorem 3.1 (label = thm:exp:inf)
Given $k\leq n\in\mathbb{N}$, $\epsilon>0$ and $m\in\mathbb{N}$ influence samples, $\mathbf{x}$}, $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ is $\epsilon$-ISDP and outputs a set of $k$ nodes which, for all $t>0$, satisfies:
$\Pr\left[I_{\mathbf{x}}(\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}}))\leq(1-e^{-1})\max_{S:|S|=k}I_{\mathbf{x}}(S)-\frac{2k^{2}n}{\epsilon m}(\ln n+t)\right]\leq ke^{-t},$
where the probability is over the randomness of the algorithm run for fixed $\mathbf{x}$}.
This theorem shows that the exponential mechanism $\mathcal{M}^{k,\epsilon}_{\mathrm{exp}}({\mathbf{x}})$ achieves $\epsilon$-ISDP while maintaining strong utility guarantees. Specifically, with high probability, the influence achieved by the private seed set is close to the optimal non-private value on the same influence samples $\mathbf{x}$}. The term (11/e)(1-1/e) reflects the classical approximation ratio for submodular maximization, while the additive term $\tfrac{2k^{2}n}{\epsilon m}(\ln n+t)$ quantifies the accuracy loss due to privacy and limited samples. In particular, the bound reveals that utility improves as the number of influence samples $m$ increases or the privacy budget $\epsilon$ becomes larger. Equipped with Proposition 2.1, we can derive the following error bound with respect to the optimal value $I_{G}$ of the influence maximization problem (proved in Section 6.4):
Theorem 3.2 (label = thm:exp:acc)
Given graph $G$ on $n\geq 2$ nodes and $k\leq n$, for any $0<\alpha\leq 1$ and $\epsilon>0$, $\mathcal{M}^{k,\epsilon}_{exp}(\cdot)$ that runs Algorithm 1 but replaces the single seed selection steps of Algorithm 2 by its exponential mechanism variant in Algorithm 3, $\mathcal{M}_{exp}^{\epsilon/k}({\mathbf{x}},S)$, is $\epsilon$-ISDP and its output infects at least $(1-1/e)\mbox{\sf OPT}-\alpha n$ nodes with high probability, using at least $m=\max\{\frac{12}{\alpha\epsilon},\frac{9}{\alpha^{2}}\}k\ln n$ influence samples in $O(knm)$ run time.
Note that if we are not concerned with privacy and set $\epsilon\gg\alpha$, then the number of samples to achieve the additive error $\alpha n$ is of the order $\alpha^{-2}k\ln n$, which improves the $k^{2}(\ln n/\alpha)/\alpha^{3}$ bound of costlyseeding. This is because in their Algorithm 1 and Theorem 2, costlyseeding collect $\rho=81klog(6nk/\epsilon)/\epsilon^{3}$ new influence samples for selecting each seed for a total of $k\left\lceil\normalcolor 81k\log({6nk}/{\epsilon})/{\epsilon^{3}}\right\rceil\normalcolor\in{O}(k^{2}\log({n}/{\epsilon})/{\epsilon^{3}})$ influence samples, whereas we observe that we can work with the same set of $m=\max\{\frac{12}{\alpha\epsilon},\frac{9}{\alpha^{2}}\}k\ln n$ influence samples to select all the $k$ seeds. Moreover, our result does not violate Theorem 3 of costlyseeding which shows that $\Omega(k^{2})$ influence samples are required to achieve the $(1-1/e)\mbox{\sf OPT}-\alpha n$ performance guarantee for all $\alpha<(1-1/e)/k$, because our algorithm also needs $\Omega(k^{2})$ samples when $\alpha<1/k$.
3.2 Local DP with Randomized Response
In the central setting, privacy protection is applied by the curator after collecting all influence samples, ensuring that the published algorithm outputs do not reveal any individual’s participation. In contrast, in the local setting, privacy is enforced during data collection itself. In contact-tracing applications, for example, it may be known that a confirmed positive case attended a particular event, but other participants may not wish to disclose their own attendance with certainty. Since individuals may not fully trust the data collector, they can protect their privacy using a randomized response mechanism—similar to those in prior local DP studies—which allows them to probabilistically deny participation. Concretely, we achieve local $\epsilon$-ISDP by perturbing the input data via randomized response: each entry of $\mathbf{x}$} is independently flipped with probability $\rho_{\epsilon}=1/(1+e^{\epsilon})$ to produce the noisy input $\tilde{{\mathbf{x}}}$, as shown in Line 3 of Algorithm 4.However, the added input noise will bias our estimate of each candidate seed’s quality; in particular, $\operatorname*{\mathbb{E}}_{\tilde{{\mathbf{x}}}}[I_{\tilde{{\mathbf{x}}}}(S|T)]$ is not necessarily equal to $I_{G}(S|T)$ for a pair of subsets $S,T\subset V$ and we no longer have the benefit of applying Proposition 2.1 directly. Rather in Algorithm 5 we propose a post-processing of the noisy input data $\tilde{{\mathbf{x}}}$ to derive an unbiased estimate of a seed set’s quality. The importance of proper post-processing steps in the analysis of DP data is highlighted in other applications, e.g., in the release of census DP data, where ignoring them can lead to biased decision-making and suboptimal results (ijcai2022p559). The added variability of the estimators due to the DP noise can also degrade our estimation performance (even if not biased), which is a second concern in local DP algorithm design and opens a venue to explore alternative post-processing methods with potentially superior bias-variance tradeoff. In the remainder of this section, we first introduce Algorithms 4 and 5 and then offer LABEL:thm:rr:acc for their privacy and utility guarantees, addressing both concerns about bias and variance of decision making with local DP influence samples.
Algorithm 4 Local $\epsilon$-ISDP seeding algorithm, $\mathcal{M}_{loc}^{k,\epsilon}({\mathbf{x}})$
1:$m$ influence samples $\mathbf{x}$}, total number of seeds $k$, and privacy budget $\epsilon>0$.
2:Set $S_{0}=\emptyset$.
3:Flip any bits of $\mathbf{x}$} with probability $\rho_{\epsilon}=\frac{1}{e^{\epsilon}+1}$ and get $\tilde{{\mathbf{x}}}$.
4:for $i\in[k]$ do
[…]
6:  $S_{i}=S_{i-1}\cup v_{i}$.
7:end for
8:Return $S_{k}$
[…]
Now, we discuss how to estimate $\Pr[z_{S}=0]$ from samples of $\tilde{z}_{S}$. The main observation is that the probability mass function of $\tilde{z}_{S}$ is a linear function of the probability mass function of $z_{S}$. Note that $S\cap x$ is the set of elements in $S$ that is also in the influence sample $x$. Hence, $S\cap\tilde{x}$ can be partitioned into (1) the subset $(S\cap\tilde{x})\cap x$ of elements of $x$ that remain in $S$ after perturbation and (2) the subset $(S\cap\tilde{x})\setminus x$ of the elements of $\tilde{x}$ that join $S$ after perturbation. The size of the first subset is a binomial random variable $Bin(|S\cap x|,1-\rho_{\epsilon})$, and the size of the second subset is $Bin(|S\setminus x|,\rho_{\epsilon})$. Therefore, given $|S|=l$ for all $a$ and $b$, the probability of $\tilde{z}_{S}=b$ conditional on $z_{S}=a$ is $\Pr[\tilde{z}_{S}=a|z_{S}=b]=\Pr[Bin(b,1-\rho_{\epsilon})+Bin(l-b,\rho_{\epsilon})=a],$ for all $a,b=0,\dots,l$ where $Bin(b,1-\rho_{\epsilon})$ and $Bin(l-b,\rho_{\epsilon})$ are independent. Furthermore, if we set the matrix $C=C_{\rho_{\epsilon},l}\in{\mathbb{R}}^{(l+1)\times(l+1)}$ as the likelihood function where $C(a,b)=C_{\rho_{\epsilon},l}(a,b)=\Pr[\tilde{z}_{S}=a|z_{S}=b]$, then we have the following:
$(\Pr[\tilde{z}_{S}=0],\ldots,\Pr[\tilde{z}_{S}=l])^{\top}=C_{\rho_{\epsilon},l}\,(\Pr[z_{S}=0],\ldots,\Pr[z_{S}=l])^{\top}.$ (2)
We can write each entry of the $C_{\rho_{\epsilon},l}$ matrix explicitly as follows:
$C_{\rho_{\epsilon},l}(a,b)=\sum_{\ell=\max(0,b-a)}^{\min(l-a,b)}\binom{b}{\ell}\binom{l-b}{a-b+\ell}\rho_{\epsilon}^{a-b+2\ell}(1-\rho_{\epsilon})^{l-a+b-2\ell}.$ (3)
To see why, consider the event that $\ell$ of the $z_{S}=b$ nodes in $S\cap x$ are flipped, leaving the $b$-$\ell$ nodes in $S\cap x\cap\tilde{x}$. To obtain $\tilde{z}_{S}=a$ nodes in $S\cap\tilde{x}$, the remaining $a-b+\ell$ nodes should come from perturbing the status of the lbl-b nodes belonging to $S\setminus x$. Note that our indices ($a$ and $b$) start at 0 instead of $1$. In the remainder of the section, we will adopt this convention.
If the matrix $C_{\rho_{\epsilon},l}$ is invertible, then we can have an unbiased estimate of $\Pr[z_{S}>0]$ using Algorithm 5, below. LABEL:prop:feasible, proved in Section 6.5, shows that $C_{\rho_{\epsilon},l}$ is invertible for $\rho_{\epsilon}>1/2$ and thus completes the specification of our randomized response algorithm.
Algorithm 5 Unbiased estimate for the size of the reachable set $J_{m}(S)$
1:$m$ perturbed influence samples $\tilde{{\mathbf{x}}}=(\tilde{x}^{1},\ldots,\tilde{x}^{m})$, their perturbation probability $\rho_{\epsilon}$, and a candidate set $S\subset V$ with size $|S|=l$.
[…]
3:Solve for ${\mathbf{f}}=(f_{0},f_{1},\dots,f_{l})\in{\mathbb{R}}^{l+1}$ so that $\tilde{{\mathbf{f}}}=C_{\rho_{\epsilon},l}{\mathbf{f}}$.
4:Return $J_{m}(S)=n(1-f_{0})$
Proposition 3.3 (label = prop:feasible, name = Feasibility of Unbiased Post-processing)
Given any positive integer $l\geq 1$ and $\rho_{\epsilon}\in[0,1/2)$, the matrix $C_{\rho_{\epsilon},l}$ is invertible.
Note that when $\rho_{\epsilon}=1/2$, $\tilde{{\mathbf{x}}}$ consists of uniform random bits and we cannot recover $\Pr[z_{S}>0]$ from perturbed influence samples. LABEL:prop:feasible shows that we can recover $\Pr[z_{S}>0]$ for any $\rho_{\epsilon}<1/2$.
Theorem 3.4 (label = thm:rr)
Given $\epsilon>0$, $m$ influence samples on a graph with $n$ nodes and integer $k<n$, Algorithm 4 is locally $\epsilon$-ISDP, and its output satisfies: $I_{G}(\mathcal{M}_{loc}^{k,\epsilon}({\mathbf{x}}))>(1-1/e)\mbox{\sf OPT}-\delta,$ with probability $1-2n^{k}\exp\left(-{m\delta^{2}}/{(2k^{2}n^{2}V_{\rho_{\epsilon},k}^{2})}\right),$ where $V_{\rho_{\epsilon},k}=O((1/2-\rho_{\epsilon})^{-k^{2}})$.
The above theorem, proved in Section 6.6, provides a tail bound on the approximation error of $\mathcal{M}_{loc}^{k,\epsilon}$ given a fixed number of influence samples. We can convert the above result to a high probability error bound by choosing $m=\Omega({k^{3}V_{\rho_{\epsilon},k}^{2}\ln n}/{\alpha^{2}})$ which for $k$ constant and $\epsilon\gg\alpha$ improves the sample complexity ($1/\alpha^{3}$) in LABEL:thm:original (complete proof in Section 6.7):
Theorem 3.5 (label = thm:rr:acc)
Given graph $G$ with $n\geq 2$ nodes and $k\leq n$, for any $0<\alpha\leq 1$ and $\epsilon>0$, $\mathcal{M}_{loc}^{k,\epsilon}$ is locally $\epsilon$-ISDP and its output infects at least $(1-1/e)\mbox{\sf OPT}-\alpha n$ nodes with high probability, using at least $m=O({k^{3}V_{\rho_{\epsilon},k}^{2}\ln n}/{\alpha^{2}})$ influence samples in $O(nk^{4}+kn^{2}m)$ run time, where $V_{\rho_{\epsilon},k}=O((1/2-\rho_{\epsilon})^{-k^{2}})$.
We conclude this section by considering two scenarios: the non-private regime $\epsilon\to\infty$ and the full-privacy regime $\epsilon\to 0$. First, in the non-private regime ($\epsilon\to\infty$), $\rho_{\epsilon}\to 0$, which causes the matrix $C_{\rho_{\epsilon},l}$ and its inverse to converge to the identity matrix, so we can take $V_{\rho_{\epsilon},k}$ to be a constant arbitrarily close to one. As a result, $m=O\left(\frac{k^{3}\ln n}{\alpha^{2}}\right)$ samples are sufficient, which is comparable to LABEL:thm:original with an additional factor of $k$, but improves the dependence on $\alpha$. We may reduce the dependence on $k$ from $k^{3}$ to a smaller order, using techniques in adaptive data analysis bassily2016algorithmic. Informally, our analysis requires that the query error in any possible seed set $S$ with $|S|\leq k$ is small enough (Lemma 6.6) so that the output of the greedy algorithm is good enough. However, we may not need such stringent conditions, because the greedy algorithm does not necessarily query all possible seed sets. Alternatively, adaptive data analysis provides a procedure to support adaptive queries from the greedy algorithm with a better error bound.  [^1] Interestingly, the procedure in bassily2016algorithmic adds Gaussian noise to each query and achieves a small error guarantee using techniques from an exponential mechanism similar to  LABEL:thm:exp:acc. Furthermore, in the non-private regime ($\epsilon\to\infty$), our exponential mechanism and the local DP seeding algorithm are identical, allowing us to use the analysis in LABEL:thm:exp:inf and derive an $O\left(\frac{k\ln n}{\alpha^{2}}\right)$ upper bound for the required number of influence samples. On the other hand, if $\epsilon\to 0$, $V_{\rho_{\epsilon},k}=O((1/2-\rho_{\epsilon})^{-k^{2}})$ in LABEL:thm:rr and LABEL:thm:rr:acc can be simplified as follows: $V_{\rho_{\epsilon},k}=O((1/2-\rho_{\epsilon})^{-k^{2}})=O((1/2-1/(e^{\epsilon}+1))^{-k^{2}})=O([2(e^{\epsilon}+1)/(e^{\epsilon}-1)]^{k^{2}})=O(\epsilon^{-k^{2}})$.
4 Simulations with Empirically-Grounded Network Data Sets
We begin by evaluating our proposed framework for data collection and intervention design using highly sensitive data on sexual activity in an HIV prevention context in Section 4.1, while additional sanity checks based on synthetic data generated from Erdős–Rényi networks are provided in Appendix 7.1.
4.1 Simulations on Empirically-Grounded Sexual Activity Networks
We construct a dynamic network of men who have sex with men (MSM) using a separable temporal exponential random graph model (STERGM), estimated from the ARTnet study (weiss2020ARTnet), a cross-sectional egocentric survey conducted in the United States (2017–2019) with 4,909 participants reporting 16,198 sexual partnerships. Our analysis focuses on San Francisco due to data availability and its representativeness for the MSM population. The STERGM estimation procedure is summarized in Appendix 7.2.
Using the estimated STERGM, we simulate sexual partnerships among 1,000 individuals over 120 weekly periods, with fixed node attributes (e.g., race, age). We aggregate every 12 weeks as one quarter. In each quarter, we randomly select 150 initial nodes and identify all individuals reachable through sexual relationships, yielding 150 influence samples per quarter and 1,500 samples in total. A second, independently generated set of influence samples is used to evaluate the seed sets output by our algorithm.
Unlike settings that assume a latent network and independent‐cascade dynamics, we directly construct influence samples to mimic contact tracing: each cascade begins with an index case and recursively includes their sexual partners. This approach reflects realistic exposure pathways without requiring full knowledge of the underlying network. To protect the privacy of individuals involved in these sensitive behavioral data, we implement both central and local privacy mechanisms: the curator adds noise to released outputs, while individuals may randomize their reported connections using randomized response.
We ran our privacy-preserving algorithm on influence samples constructed from ARTnet study data. In this case, the chosen seeds are high-risk nodes whose protection provides the most overall benefit in terms of preventing large disease cascades. In the simulation setup of Figure 1, we consider $k=18$ and $k=20$ seeds, the number of influence samples ranges from zero to $1500$, $m\in\{0,500,1000,1500\}$. The case $m=0$ corresponds to the absence of input information where the choice of 18 or 20 seeds will be uniformly random. With no input information at $m=0$, there are no privacy risks to consider.
Figure 1: Performance of our privacy-preserving seeding algorithms—the exponential-mechanism and randomized-response methods—on the MSM dataset. We report the expected influence $I_{G}(\cdot)$ across privacy budgets ($\epsilon$) as the number of influence samples ($m$) increases, and compare against two non-private baselines: influence-sample seeding and the deterministic greedy algorithm with complete network information. Each point represents the mean over 1,000 trials obtained from 50 independent collections of $m$ influence samples, each evaluated 20 times. Error bars show $95\%$ confidence intervals (mostly too small to be visible). The expected spread size for each seed set is estimated using an additional 1,000 influence samples.
In the Figure 1, we see that the exponential mechanism outperforms the randomized response mechanism as the latter provides stronger (local) privacy protections. We also observe that low values of privacy budget $\epsilon$ lead to earlier saturation in lower expected spread sizes. In those cases, our maximum number of influence samples is not enough to counteract the random noise introduced by privacy mechanisms, so the value of network information diminishes with the decreasing privacy budge – cf. costlyseeding. Increasing the privacy budget alleviates this effect, as we see a faster increase in the expected spread size with increasing $m$ at higher values of $\epsilon$. In fact, the performance of both mechanisms approaches the original (“non-private”) influence sample greedy seeding algorithm (Algorithm 1), as $\epsilon\to\infty$. For example, in the randomized response mechanism (Algorithm 4), as the privacy budget increases, the flip probability of the entries in the influence samples $x\in\{0,1\}^{n\times m}$ (Algorithm 4, line 3) decreases and the matrix $C_{\rho_{\epsilon},l}$ in Equation 3 approaches the identity matrix, so the performance is similar to greedy non-private seeding on influence samples in Algorithm 1. For the exponential mechanism of Section 3.1, at each step the mechanism chooses the seeds probabilistically with probabilities proportional to the exponential of their marginal contribution to the spread size computed on the influence samples. In this case, increasing the privacy budget ($\epsilon\to\infty$) increases the probability of choosing the seed with the largest marginal contribution. Figure 1 also shows the deterministic baseline for greedy with complete network information to be compared with the exponential mechanism and randomized response algorithms at different $\epsilon$ values with increasing $m$. As we gain better information about the network by increasing the number of influence samples $(m\to\infty)$ and the privacy budget $(\epsilon\to\infty)$, the performance of both algorithms will converge to the deterministic greedy algorithm with complete network information.
5 Discussions, managerial insights and future directions
Comparison between node, edge, and influence-sample DP.
We introduce a new notion of differential privacy—influence-sample differential privacy (ISDP)—which defines adjacency based on differences in observed diffusion samples rather than changes to the underlying network structure. This formulation is particularly suited to applications where the network is unknown or cannot be reliably reconstructed due to privacy constraints, such as HIV prevention, election polling, or information diffusion on social media. In contrast to node- or edge-level DP, ISDP does not assume access to the full network. When the network is latent, modifying a single node or edge can alter the entire diffusion process, thereby changing many entries in all observed cascades. Satisfying node or edge DP in such cases would require adding excessive noise, effectively destroying utility. ISDP circumvents this limitation by defining adjacency at the level of individual cascade participation—whether a person appears in a given diffusion sample—offering a meaningful privacy guarantee even without network knowledge.
This distinction highlights that different definitions of adjacency can lead to fundamentally different trade-offs between privacy and utility (Propositions 2.5 and 2.6 versus LABEL:thm:exp:acc and LABEL:thm:rr:acc). Our goal is not to claim ISDP as superior to node or edge DP, but to emphasize that the appropriate notion of adjacency should reflect the structure and semantics of the data. In networked settings, individuals’ private data—be it census, medical, or behavioral records—are inherently correlated through network ties. Those supporting node or edge DP may argue that adjacency should depend on the network structure rather than the traditional Hamming distance between database entries (rezaei2019privacy). Conceptually, ISDP protects whether an individual is directly or indirectly connected to high-risk individuals as inferred from observed cascades, rather than the presence of network edges themselves. By focusing on influence samples, ISDP captures the most relevant privacy risk in sensitive contexts while maintaining practical utility for data-driven interventions.
Efficiency and robustness.
Our algorithms have polynomial running times in the size of the graph, with a time complexity of $O(knm)$ and $O(nk^{4}+kn^{2}m)$ for $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ and $\mathcal{M}_{loc}^{k,\epsilon}({\mathbf{x}})$, respectively, for any constant privacy budget $\epsilon>0$. To improve running times, one may consider improving the iterative sampler for the exponential mechanism in Algorithm 3, or designing a better post-processing step for Line 3 in Algorithm 5. On the other hand, our ISDP randomized response algorithm $\mathcal{M}_{loc}^{k,\epsilon}({\mathbf{x}})$ is automatically robust against independent random noise in influence samples. In contrast, it is unclear how well our ISDP exponential mechanism $\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}})$ performs in the presence of noise.
Conclusions and future directions.
By focusing on the utility of influence samples for influence maximization, our approach ensures that each user’s participation in any cascade remains private—either locally or centrally—while preserving the statistical validity of the cascades for influence estimation. We formally protect whether an individual appears in a given influence sample, granting participants plausible deniability regarding their involvement in diffusion events. Our framework applies broadly to settings such as email communication, social messaging, and ad propagation, where actions are observed over time but the underlying network is unknown or infeasible to collect in full. In future work, we plan to extend ISDP to temporally evolving data, investigating how sequential data collection affects both privacy guarantees and intervention effectiveness. Such extensions may lead to new, consequentialist notions of privacy that remain robust under varying sampling distributions and randomization noise (kifer2014pufferfish, liu2020learning).
Code and Data Availability
The MSM sexual behavior cascade data are derived from the ARTnet study (https://github.com/EpiModel/ARTnet). Due to confidentiality agreements, the dataset is not publicly available; access requires a memorandum of understanding (MOU) and approval from the study’s Principal Investigator, Samuel Jenness (Emory University). The analysis codes used to produce the results and figures are available at https://github.com/aminrahimian/dp-inf-max.
{APPENDICES}
6 Proofs and Supplementary Results
6.1 Proof of Proposition 2.5
For the graph with maximum degree n1n-1, consider the isolated graph $G^{\prime}=(V,p^{\prime})$ of $n$ nodes where $p^{\prime}_{e}=0$ for all possible pairs $e$. Given any $\epsilon$ node-DP mechanism $M$, there exists a node $v_{M}\in V$ such that $\Pr[M^{k=1}(G^{\prime})=v_{M}]\leq 1/n,$ according to the pigeonhole principle. Then we define a star graph $G=(V,p)$ of $n$ nodes with center $v_{M}$ where $p_{(v_{M},v)}=1/2$ for all $v\in V$ and $v\neq v_{M}$ and $p_{(u,v)}=0$ if $u,v\neq v_{M}$.
First, the center node can in expectation infect (n1)/2+1(n-1)/2+1 of the nodes in $G$, therefore, $\mbox{\sf OPT}\geq I_{G}(\{v_{M}\})=(n+1)/2$. On the other hand, because $M^{k}$ is $\epsilon$ node-DP and $G$ and $G^{\prime}$ are node-adjacent, $\Pr[M^{k=1}(G)=\{v_{M}\}]\leq e^{\epsilon}\Pr[M^{k=1}(G^{\prime})=\{v_{M}\}]\leq{e^{\epsilon}}/{n}$. Therefore,
$\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]=$ $\operatorname*{\mathbb{E}}[I_{G}({\mathbf{y}})|{\mathbf{y}}\neq\{v_{M}\}]\Pr[M^{k=1}(G)\neq\{v_{M}\}]+\operatorname*{\mathbb{E}}[I_{G}(\{v_{M}\})]\Pr[M^{k=1}(G)=\{v_{M}\}]$
$\leq$ $(1+1/2+(n-2)/4)+n\cdot{e^{\epsilon}}/{n}=1+e^{\epsilon}+n/4,$
For graphs with maximum degree $2\leq d<n-1$, we construct $G^{\prime}=(V,p^{\prime})$ consisting of $l$ disjoint (d1)(d\!-\!1)-regular trees $S_{1},\ldots,S_{l}$, each of size about $n/l$ (a complete d1d-1-regular tree truncated to the largest depth not exceeding $n/l$, with the last level partially filled if necessary). For each edge $e$, set $p_{e}^{\prime}=1$ if $e$ belongs to some $S_{i}$ and $p_{e}^{\prime}=0$ otherwise. By the pigeonhole principle, there exist $a$ trees $S_{i_{1}},\ldots,S_{i_{a}}$ such that $\Pr[M^{k=1}(G^{\prime})\subset S_{i_{1}}\cup\cdots\cup S_{i_{a}}]\leq a/l$.
We then obtain $G=(V,p)$ by adding a new node $v$ and connecting it to the roots of $S_{i_{1}},\ldots,S_{i_{a}}$. Any $u\in S_{i_{1}}\cup\cdots\cup S_{i_{a}}$ now reaches all nodes in these $a$ trees, giving $I_{G}(u)=an/l+1$, while any $u\notin S_{i_{1}}\cup\cdots\cup S_{i_{a}}$ reaches only its own tree, so $I_{G}(u)=n/l$.
Because $M$ is $\epsilon$-node DP and $G$ and $G^{\prime}$ differ by only one node, we have $\Pr[M^{k=1}(G)\subset S_{i_{1}}\cup\cdots\cup S_{i_{a}}]\leq ae^{\epsilon}/l$. Hence
$\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\leq\frac{ae^{\epsilon}}{l}\Bigl(\frac{an}{l}+1\Bigr)+\Bigl(1-\frac{ae^{\epsilon}}{l}\Bigr)\frac{n}{l}=\Bigl(\frac{1}{a}+\frac{(a-1)e^{\epsilon}}{l}\Bigr)\frac{an}{l}+\frac{ae^{\epsilon}}{l}.$
Comparing with $\mbox{\sf OPT}=an/l$ gives
$(1-1/e)\mbox{\sf OPT}-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\geq\Bigl(\frac{a-1}{a}-\frac{1}{e}-\frac{(a-1)e^{\epsilon}}{l}\Bigr)\frac{an}{l}-\frac{ae^{\epsilon}}{l}.$
We choose $l$ so that $\frac{a-1}{a}-\frac{1}{e}-\frac{(a-1)e^{\epsilon}}{l}=\frac{1}{10}$. Since the coefficient of $n$ in the resulting lower bound is increasing in $a$ and $a\leq d$, setting $a=d$ gives
$(1-1/e)\mbox{\sf OPT}-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M(G)}[I_{G}({\mathbf{y}})]\geq\frac{1}{10e^{\epsilon}}\Bigl(1-\frac{d}{d-1}\Bigr)\Bigl(\frac{1}{e}+\frac{1}{10}\Bigr)n-1+\frac{d}{d-1}\Bigl(\frac{1}{e}+\frac{1}{10}\Bigr).$
which completes the proof. $\Box$
6.2 Proof of Proposition 2.6
Consider $G^{\prime}=(V,p^{\prime})$ on $n$ nodes formed by $l$ disjoint d1d-1-regular trees $S_{1},\dots,S_{l}$, each of size about $n/l$ (with the last level truncated if needed). For each edge $e$, set $p^{\prime}_{e}=1$ if $e$ lies in some $S_{i}$ and $p^{\prime}_{e}=0$ otherwise. As in Proposition 2.5, for any $\epsilon$-edge DP mechanism $M$ there exist two trees $S_{i},S_{j}$ such that $\Pr[M^{k=1}(G^{\prime})\subset S_{i}\cup S_{j}]\leq 2/l$.
Define $G=(V,p)$ by adding a single bridge edge $e^{*}$ between the roots of $S_{i}$ and $S_{j}$, with $p_{e^{*}}=1$ and $p_{e}=p^{\prime}_{e}$ for all $e\neq e^{*}$. Then any $v\in S_{i}\cup S_{j}$ reaches $2n/l$ nodes in $G$, so $I_{G}(v)=2n/l$, whereas any $v\notin S_{i}\cup S_{j}$ reaches only its own tree, so $I_{G}(v)=n/l$. Because $M$ is $\epsilon$-edge DP and $G$ and $G^{\prime}$ differ in a single edge, we have $\Pr[M^{k=1}(G)\subset S_{i}\cup S_{j}]\leq 2e^{\epsilon}/l$. Hence
$\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\leq\frac{2e^{\epsilon}}{l}\cdot\frac{2n}{l}+\Bigl(1-\frac{2e^{\epsilon}}{l}\Bigr)\frac{n}{l}=\Bigl(\frac{1}{2}+\frac{e^{\epsilon}}{l}\Bigr)\frac{2n}{l},$
and therefore
$(1-1/e)\mbox{\sf OPT}-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\geq\Bigl(\frac{1}{2}-\frac{1}{e}-\frac{e^{\epsilon}}{l}\Bigr)\frac{2n}{l}.$
We choose $l$ so that $\frac{1}{2}-\frac{1}{e}-\frac{e^{\epsilon}}{l}=\frac{1}{10}$, which yields
$(1-1/e)\mbox{\sf OPT}-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M(G)}[I_{G}({\mathbf{y}})]\geq\frac{1}{10e^{\epsilon}}\Bigl(\frac{4}{5}-\frac{2}{e}\Bigr)n-\frac{4}{5}+\frac{2}{e},$
which completes the proof.
$\Box$
One may wonder if the lower bounds hold for approximate node- and edge-DP. Formally, an algorithm $M$ is $(\epsilon,\delta)$-node- (or edge-) DP if for all outputs $\mathbf{y}$} and node- or edge-adjacent graphs $G,G^{\prime}$, $\Pr[M(G)={\mathbf{y}}]\leq e^{\epsilon}\Pr[M(G^{\prime})={\mathbf{y}}]+\delta$. We often consider $\delta$ to be negligibly small, so the algorithms are approximately $\epsilon$-DP. It is not hard to see that our proof for Propositions 2.5 and 2.6 still holds for the approximated DP. For example,
Proposition 6.1
For all $n\geq 1$, $\epsilon>0$ and $\delta<1/6$, there exists $\alpha_{\epsilon}\geq\frac{1}{10000e^{\epsilon}}$ so that any $(\epsilon,\delta)$-edge-DP seeding algorithm $M^{k}(G)$, there exists graph $G$ on $n$ nodes so that for $k=1$: $(1-1/e)\mbox{\sf OPT}-\operatorname*{\mathbb{E}}_{{\mathbf{y}}\sim M^{k=1}(G)}[I_{G}({\mathbf{y}})]\geq\alpha_{\epsilon}n.$
The proof is mostly identical to Proposition 2.6, except that when $M$ is $(\epsilon,\delta)$-DP, we only have $\Pr[M^{k=1}(G)\subset S_{i}\cup S_{j}]\leq{2e^{\epsilon}}/{l}+\delta$ . Then the rest of the proof follows.
6.3 Proof of LABEL:thm:exp:inf
Because $\mathbf{x}$} and $\mathbf{x}$$^{\prime}$ are adjacent if those matrices are different at one position, the sensitivity of the scoring function $I_{\mathbf{x}}(v|S)$ for any $S$ is
$\max_{v\in[n]}\max_{{\mathbf{x}}\underset{is}{\sim}{\mathbf{x}}^{\prime}}|I_{\mathbf{x}}(v|S)-I_{{\mathbf{x}}^{\prime}}(v|S)|$ $\leq\frac{n}{m}\max_{v\in[n]}\max_{{\mathbf{x}}\underset{is}{\sim}{\mathbf{x}}^{\prime}}\left|\sum^{m}_{t=1}x^{t}_{v}-(x^{\prime})^{t}_{v}\right|=\frac{n}{m},$
and the size of the output space is $n$. Therefore for all $S$, Algorithm 3 is an $\epsilon$-DP exponential mechanism.
Privacy: Let the output be $(v_{1},\dots,v_{k})$. Because for any fixed $S$, Algorithm 3 is an exponential mechanism, we have
$\ln\frac{\Pr[\mathcal{M}_{exp}^{\epsilon/k}({\mathbf{x}},S)=v_{i}]}{\Pr[\mathcal{M}_{exp}^{\epsilon/k}({\mathbf{x}}^{\prime},S)=v_{i}]}\leq\frac{\epsilon}{k}\text{ for all set }S\text{ and adjacent }{\mathbf{x}}\underset{is}{\sim}{\mathbf{x}}^{\prime}.$
Then we can write out the probability ratio of $\mathcal{M}_{exp}^{\epsilon}$ outputting the same seed set $v_{1},\dots,v_{k}$ on two adjacent datasets $\mathbf{x}$} and $\mathbf{x}$$^{\prime}$,
$\ln\frac{\Pr[\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}})=(v_{1},\dots,v_{k})]}{\Pr[\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}}^{\prime})=(v_{1},\dots,v_{k})]}=$ $\sum_{i}\ln\frac{\Pr[\mathcal{M}_{exp}^{\epsilon/k}({\mathbf{x}},S_{i-1})=v_{i}\mid S_{i-1}=(v_{1},\dots,v_{i-1})]}{\Pr[\mathcal{M}_{exp}^{\epsilon/k}({\mathbf{x}}^{\prime},S_{i-1})=v_{i}\mid S_{i-1}=(v_{1},\dots,v_{i-1})]}$
$\leq$ $\sum_{i}\epsilon/k=\epsilon.$
Utility guarantee: Given $\mathbf{x}$}, we show that the output of $\mathcal{M}_{exp}^{\epsilon}$ is close to the optimal. First note that once $\mathbf{x}$} is fixed, Algorithm 1 is the classical greedy algorithm for the objective function $I_{\mathbf{x}}$ that is submodular. Second, the error comes from the soft max of Algorithm 3: at each $i\in[k]$, instead of choosing the seed $v^{*}_{i}$ to maximize the score, Algorithm 3 chooses a random seed $v_{i}$ according to the probabilities specified in its line 5.
We set the optimal set for $I_{\mathbf{x}}$ as $S^{*}:=\{v^{*}_{1},\dots,v^{*}_{k}\}$ so that $I_{\mathbf{x}}(S^{*})=\max_{S:|S|\leq k}I_{\mathbf{x}}(S)$. Given a seed set $v_{1},\dots,v_{k}$, for all $i=0,\dots,k$, let $S_{i}:=\{v_{j}:1\leq j\leq i\}$ and define the remainder of $S_{i}$ as: $R_{i}:=I_{\mathbf{x}}(S^{*})-I_{\mathbf{x}}(S_{i})=\frac{n}{m}(|C_{\mathbf{x}}(S^{*})|-|C_{\mathbf{x}}(S_{i})|).$ Because $R_{i+1}=R_{i}-(I_{\mathbf{x}}(S_{i+1})-I_{\mathbf{x}}(S_{i}))$, to obtain an upper bound $R_{i+1}$, it suffices to find a lower bound on $I_{\mathbf{x}}(S_{i+1})-I_{\mathbf{x}}(S_{i})$. Then, because $C_{\mathbf{x}}(S_{i})\subseteq C_{\mathbf{x}}(S_{i+1})$, $I_{\mathbf{x}}(S_{i+1})-I_{\mathbf{x}}(S_{i})=\frac{n}{m}(|C_{\mathbf{x}}(S_{i+1})|-|C_{\mathbf{x}}(S_{i})|)=\frac{n}{m}|C_{\mathbf{x}}(S_{i+1})\setminus C_{\mathbf{x}}(S_{i})|=I_{\mathbf{x}}(v_{i+1}|S_{i})$. Thus, we have $R_{i+1}=R_{i}-I_{\mathbf{x}}(v_{i+1}|S_{i})$. On the other hand, we show a lower bound of $I_{\mathbf{x}}(v_{i+1}|S_{i})$. First, we want to apply LABEL:thm:exp:utility. Since the sensitivity is $n/m$, the size of the output space is $n$ for $\mathcal{M}_{exp}^{\epsilon/k}$, given $t>0$ we can set $\beta=\frac{2nk}{\epsilon m}(\ln n+t)$, and by LABEL:thm:exp:utility and the union bound we have for all $i=1,\dots,k$: $I_{\mathbf{x}}(v_{i+1}|S_{i})\geq\max_{w\in[n]}I_{\mathbf{x}}(w|S_{i})-\beta,$ with probability $k\exp(-t)$. Additionally, we have:
$\max_{w}I_{\mathbf{x}}(w|S_{i})\geq$ $\frac{1}{k}\sum_{j=1}^{k}I_{\mathbf{x}}(v_{j}^{*}|S_{i})=\frac{1}{k}\sum_{j}\frac{n}{m}|C_{\mathbf{x}}(v_{j}^{*}\cup S_{i})\setminus C_{\mathbf{x}}(S_{i})|$
$\geq$ $\frac{1}{k}\frac{n}{m}\left|\cup_{j=1}^{k}C_{\mathbf{x}}(v_{j}^{*}\cup S_{i})\setminus C_{\mathbf{x}}(S_{i})\right|=\frac{1}{k}I_{\mathbf{x}}(S^{*}|S_{i})=\frac{1}{k}R_{i}.$
Therefore, we can combine these two inequalities and have $R_{i+1}=R_{i}-I_{\mathbf{x}}(v_{i+1}|S_{i})\leq R_{i}+\beta-\max_{w}I_{\mathbf{x}}(w|S_{i})\leq(1-1/k)R_{i}+\beta$. Because $R_{k}=I_{\mathbf{x}}(S^{*})-I_{\mathbf{x}}(S_{k})$ and $R_{0}=I_{\mathbf{x}}(S^{*})$, the error is bounded by
$R_{k}\leq(1-1/k)^{k}R_{0}+\sum_{i=1}^{k}(1-1/k)^{k-i}\beta\leq\frac{1}{e}\max_{S:|S|=k}I_{\mathbf{x}}(S)+k\beta,$
where $k\beta=\frac{2nk^{2}}{\epsilon m}(\ln n+t)$. $\Box$
Lemma 6.2 (label = thm:exp:utility, name = Exponential mechanism for selecting a greedy seed)
Given $k=1$, $\epsilon>0$ and $m\in\mathbb{N}$ influence samples, $\mathbf{x}$}, $\mathcal{M}^{k=1,\epsilon}_{exp}({\mathbf{x}})$ is $\epsilon$-ISDP and outputs a single greedy node which, for all $t>0$, satisfies:
$\Pr\left[I_{\mathbf{x}}(\mathcal{M}^{k,\epsilon}_{exp}({\mathbf{x}}))\leq\max_{S:|S|=1}I_{\mathbf{x}}(S)-\frac{2n}{\epsilon m}(\ln n+t)\right]\leq e^{-t},$
where the probability is over the randomness of the algorithm run for fixed $\mathbf{x}$}.
This is an immediate application of Theorem 3.11 of dwork2014algorithmic. Here the utility function is $u({\mathbf{x}},v)=I_{\mathbf{x}}(v\mid S)=\tfrac{n}{m}\,|C_{\mathbf{x}}(S\cup\{v\})\setminus C_{\mathbf{x}}(S)|$, whose sensitivity is $\Delta(u)=\tfrac{n}{m}$. Plugging this sensitivity into Theorem 3.11 yields the claim.
6.4 Proof of LABEL:thm:exp:acc
Because $m\geq{9}k(\ln n)/{\alpha^{2}}$, we can use Proposition 2.1, together with a union bound on all subsets of size at most $k$, to guarantee that $|I_{\mathbf{x}}(S)-I_{G}(S)|\leq{\alpha n}/{3},$ for all $S$ with $|S|\leq k$ with probability at least $1-2n^{k}\exp(-{2\alpha^{2}}m/{9})\geq 1-2n^{k}\exp(-2k\ln n)=1-2n^{-k}$. Therefore,
$I_{G}(\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}}))\geq I_{\mathbf{x}}(\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}}))-\frac{\alpha n}{3},$ (4)
$\mbox{\sf OPT}=\max_{S:|S|=k}I_{G}(S)\leq\max_{S:|S|=k}I_{\mathbf{x}}(S)+\frac{\alpha n}{3}.$ (5)
On the other hand, by LABEL:thm:exp:inf, $\mathcal{M}_{exp}^{\epsilon}(\cdot)$ is $\epsilon$-ISDP. Moreover, if we set $t=\ln n$, because $m\geq\frac{12}{\alpha\epsilon}k\ln n$, we have:
$I_{\mathbf{x}}(\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}}))$ $\geq(1-1/e)\max_{S:|S|=k}I_{\mathbf{x}}(S)-\frac{2kn}{\epsilon m}\cdot 2\ln n$
$\geq(1-1/e)\max_{S:|S|=k}I_{\mathbf{x}}(S)-\frac{\alpha n}{3},$
and
$I_{\mathbf{x}}(\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}}))\geq(1-1/e)\max_{S:|S|=k}I_{\mathbf{x}}(S)-\frac{\alpha n}{3},$ (6)
with probability at least $1-\exp(-\ln n)=1-1/n$. Combining Equations 4, 5 and 6, with probability at least $1-2n^{-k}-1/n=1-O(1/n)$ we have:
$I_{G}(\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}}))\geq$ $I_{\mathbf{x}}(\mathcal{M}_{exp}^{\epsilon}({\mathbf{x}}))-\frac{\alpha n}{3}$ (by Equation 4)
$\geq$ $(1-1/e)\max_{S:|S|=k}I_{\mathbf{x}}(S)-\frac{2\alpha n}{3}$ (by Equation 6)
$\geq$ $(1-1/e)\mbox{\sf OPT}-\alpha n,$ (by Equation 5)
which completes the proof.
Run time: Note that the time complexity of computing $I_{\mathbf{x}}(v|S)$ is $m$, which gives a $O(knm)$ bound on the time complexity of the exponential mechanism seeding algorithm, iterating the single seed selection — Algorithm 3kk times. $\Box$
6.5 Proof of LABEL:prop:feasible
Suppose $C=C_{\rho,l}$ is singular for some $\rho$ and there exist two distinct vectors ${\mathbf{v}},{\mathbf{u}}\in{\mathbb{R}}^{l+1}$ such that $C{\mathbf{v}}=C{\mathbf{u}}$. In this proof, we will use $a$ and $b$ as indices varying from 0 to $l$. We now assume that ${\mathbf{v}}=(v_{0},\dots,v_{l})$ and ${\mathbf{u}}=(u_{0},\dots,u_{l})$ are two distinct distributions such that $v_{b},u_{b}\geq 0$ for all $b$ and $\sum_{b}v_{b}=\sum_{b}u_{b}=1$. We will relax this assumption later. Because $C$ is a likelihood function where $C(a,b)=\Pr[\tilde{z}_{S}=a|z_{S}=b]=\Pr[Bin(b,1-\rho)+Bin(l-b,\rho)=a]$ the product $Cv\in{\mathbb{R}}^{l+1}$ is also a distribution. Now we compute the moment-generating function of a random variable $y$ with distribution $Cv$. Given $0\leq a\leq l$, let $y_{b}:=Bin(b,1-\rho)+Bin(l-b,\rho)$ be the random variable of the sum of two independent binomial random variables. Then the moment-generating function of $y_{b}$ is $\operatorname*{\mathbb{E}}[e^{ty_{b}}]=(\rho+(1-\rho)e^{t})^{b}(1-\rho+\rho e^{t})^{l-b}$. Moreover, because the random variable $y$ with distribution $Cv$ is the mixture of $y_{b}$ with weight $v_{b}$, $y=\sum_{b=0}^{l}v_{b}y_{b}$, the moment generating function of $y$ is
$\textstyle\operatorname*{\mathbb{E}}_{y\sim C{\mathbf{v}}}[e^{ty}]=$ $\sum_{b=0}^{l}v_{b}\operatorname*{\mathbb{E}}[e^{ty_{b}}]=\sum_{b=0}^{l}v_{b}(1-\rho+\rho e^{t})^{b}(\rho+(1-\rho)e^{t})^{l-b}$
$=$ $(\rho+(1-\rho)e^{t})^{l}\sum_{b=0}^{l}v_{b}\left(\frac{1-\rho+\rho e^{t}}{\rho+(1-\rho)e^{t}}\right)^{b},$
for all $t$. Similarly, the moment generating function of a random variable with distribution $C{\mathbf{u}}$ is
$\operatorname*{\mathbb{E}}_{y\sim C{\mathbf{u}}}[e^{ty}]=(\rho+(1-\rho)e^{t})^{l}\sum_{b=0}^{l}u_{b}\left(\frac{1-\rho+\rho e^{t}}{\rho+(1-\rho)e^{t}}\right)^{b}.$
Because $C{\mathbf{u}}=C{\mathbf{v}}$, the moment generating functions are equal, $\operatorname*{\mathbb{E}}_{y\sim C{\mathbf{v}}}[e^{ty}]=\operatorname*{\mathbb{E}}_{y\sim C{\mathbf{u}}}[e^{ty}]$ for all $t$. Since $(1-\rho+\rho e^{t})^{l}>0$ for all $t$, we have
$\sum_{b=0}^{l}(v_{b}-u_{b})\left(\frac{\rho+(1-\rho)e^{t}}{1-\rho+\rho e^{t}}\right)^{b}=0,\forall t.$
Because $\rho\neq 1/2$, $\frac{\rho+(1-\rho)e^{t}}{1-\rho+\rho e^{t}}\neq\frac{\rho+(1-\rho)e^{t^{\prime}}}{1-\rho+\rho e^{t^{\prime}}}$ for all $t\neq t^{\prime}$, so the degree $l$ polynomial $\sum_{b=0}^{l}(v_{b}-u_{b})x^{b}$ has more than $l$ zeros. By the fundamental theorem of algebra, ${\mathbf{v}}={\mathbf{u}}$.
Finally, suppose that there exists ${\mathbf{v}}\neq{\mathbf{u}}$ with $C{\mathbf{v}}=C{\mathbf{u}}$ but $\mathbf{v}$} is not a distribution. We show that there exist two distributions ${\mathbf{v}}^{\prime}\neq{\mathbf{u}}^{\prime}$ with $C{\mathbf{v}}^{\prime}=C{\mathbf{u}}^{\prime}$ which completes the proof. Let $\beta:=\min_{b}\{v_{b},u_{b}\}$. Because each row of $C$ is a distribution and $\mathbf{1}^{\top}C=\mathbf{1}^{\top}$, we can set $\alpha:=\sum_{b}v_{b}=\mathbf{1}^{\top}{\mathbf{v}}=\mathbf{1}^{\top}C{\mathbf{v}}=\mathbf{1}^{\top}C{\mathbf{u}}=\sum_{b}u_{b}$. We can see that the sum of $l+1$ terms should be greater than the minimum times $l+1$, $\alpha-(l+1)\beta>0$; otherwise, $v_{b}=u_{b}=\beta$ for all $b$ and ${\mathbf{v}}={\mathbf{u}}$. Then we define ${\mathbf{v}}^{\prime},{\mathbf{u}}^{\prime}\in{\mathbb{R}}^{l+1}$ so that $v^{\prime}_{b}=\frac{v_{b}-\beta}{\alpha-(l+1)\beta}$ and $u^{\prime}_{b}=\frac{u_{b}-\beta}{\alpha-(l+1)\beta}$ for all $b$. We can see that $\mathbf{v}$$^{\prime}$ and $\mathbf{u}$$^{\prime}$ are two distinct distributions and that $C{\mathbf{v}}^{\prime}=\frac{1}{\alpha-(l+1)\beta}C{\mathbf{v}}-\frac{1}{\alpha-(l+1)\beta}C\mathbf{1}=\frac{1}{\alpha-(l+1)\beta}C{\mathbf{u}}-\frac{1}{\alpha-(l+1)\beta}C\mathbf{1}=C{\mathbf{u}}^{\prime},$ completes the proof. $\Box$
6.6 Proof of LABEL:thm:rr
Because all bits of the influence samples are flipped with probability $\rho_{\epsilon}$, the perturbed $\tilde{{\mathbf{x}}}$ is $\epsilon$ locally DP. By the post-processing immunity of DP, the output of Algorithm 4 is $\epsilon$-locally DP.
For the utility, by Lemmas 6.4 and 6.6 and a union bound, for any $\delta^{\prime}>0$ and $S$ with $|S|\leq k$, the estimate $J_{m}$ is close to the influence function $I_{G}$
$|J_{m}(S)-I_{G}(S)|\leq\delta^{\prime},$ (7)
with probability $1-2n^{k}\exp(-\frac{2m\delta^{\prime 2}}{n^{2}V^{2}_{\rho_{\epsilon},k}})$ over the randomness of the perturbation.
[…]
$I_{G}(S\cup u)\geq$ $J_{m}(S\cup u)-\delta^{\prime}\geq J_{m}(S\cup u^{*})-\delta^{\prime}$
$\geq$ $I_{G}(S\cup u^{*})-2\delta^{\prime}=\max_{v}I_{G}(S\cup v)-2\delta^{\prime}.$
[…]
$I_{G}(\mathcal{M}_{loc}^{k,\epsilon}({\mathbf{x}}))$ $>(1-1/e)\mbox{\sf OPT}-\sum_{i=1}^{k}(1-1/k)^{k-i}\delta_{i}$
$\geq(1-1/e)\mbox{\sf OPT}-2k\delta^{\prime},$
with probability $1-2n^{k}\exp(-\frac{2m\delta^{\prime 2}}{n^{2}V^{2}_{\rho_{\epsilon},k}})$. We complete the proof by setting $\delta^{\prime}=\frac{\delta}{2k}$. $\Box$
Lemma 6.3
Suppose $A:2^{V}\to V$ is an iterative single seed selection algorithm with a sequence of errors $\delta_{0},\delta_{1},\dots,\delta_{k-1}>0$ so that for all $S\subseteq[n]$ with $|S|=i<k$, $I_{G}(S\cup A(S))>\max_{v}I_{G}(S\cup v)-\delta_{i}$. If $S_{i}=S_{i-1}\cup A(S_{i-1})$ for all $0<i\leq k$ and $S_{0}=\emptyset$, then $I_{G}(S_{k})>(1-1/e)\mbox{\sf OPT}-\sum_{i=1}^{k}(1-1/k)^{k-i}\delta_{i}.$
Proof of Lemma 6.3: We set the optimal $k$ seeds for $I_{G}$ as $v_{1}^{*}$, $\dots,v_{k}^{*}$ so that $I_{G}(\{v_{1}^{*},\dots,v_{k}^{*}\})=\max_{S:|S|\leq k}I_{G}(S)=\mbox{\sf OPT}$ and let $v_{1},\dots,v_{k}$ be the outputs of algorithm $A$ so that $S_{0}=\emptyset$ and $S_{i}=\{v_{1},\dots,v_{i}\}=S_{i-1}\cup A(S_{i-1})$ for all $1\leq i\leq k$. We define the remainder of $S_{i}$ as $R_{i}=I_{G}(S^{*})-I_{G}(S_{i})$ which is the gap between the influence of the optimal seed set $S^{*}$ and $S_{i}$. Note that $R_{0}=\mbox{\sf OPT}$. We want prove that
$R_{i+1}\leq(1-1/k)R_{i}+\delta_{i}\text{, for all }i=0,\dots,k-1.$ (8)
That will complete the proof, because $R_{k}\leq(1-1/k)^{k}R_{0}+\sum_{i}(1-1/k)^{k-i}\delta_{i}\leq\frac{1}{e}\mbox{\sf OPT}+\sum_{i}(1-1/k)^{k-i}\delta_{i}$ and $R_{k}=\mbox{\sf OPT}-I_{G}(S_{k})$, which together imply $\mbox{\sf OPT}-I_{G}(S_{k})\leq\frac{1}{e}\mbox{\sf OPT}+\sum_{i}(1-1/k)^{k-i}\delta_{i}$.
Now we prove (8). Similarly to the proof of LABEL:thm:exp:inf, first by definition, we have $R_{i+1}=R_{i}-(I_{G}(S_{i+1})-I_{G}(S_{i}))$. Second, because $S_{i}\subset S_{i+1}$, $I_{G}(S_{i+1})-I_{G}(S_{i})=I_{G}(v_{i+1}|S_{i})$, we have: $R_{i+1}=R_{i}-I_{G}(v_{i+1}|S_{i}).$ To lower bound $I_{G}(v_{i+1}|S_{i})$, by the condition of the lemma, we have:
$I_{G}(v_{i+1}\mid S_{i})=I_{G}(S_{i}\cup v_{i+1})-I_{G}(S_{i})\geq\max_{v}I_{G}(S_{i}\cup v)-\delta_{i}-I_{G}(S_{i})=\max_{v}I_{G}(v\mid S_{i})-\delta_{i}.$
Finally, because $I_{G}(\cdot|S_{i})$ is submodular, $I_{G}(S^{*}|S_{i})\leq\sum_{j=1}^{k}I_{G}(v_{j}^{*}|S_{i})\leq k\max_{v}I_{G}(v|S_{i})$. By definition, $R_{i}\leq I_{G}(S^{*}|S_{i})$, and we have $\max_{v}I_{G}(v|S_{i})\geq\frac{1}{k}R_{i}.$ We can combine the above three inequalities and prove (8), $R_{i+1}=R_{i}-I_{G}(v_{i+1}|S_{i})\leq R_{i}-\max_{v}I_{G}(v|S_{i})+\delta_{i}\leq(1-1/k)R_{i}+\delta_{i}$. $\Box$
By LABEL:prop:feasible, we know that $C_{\rho,l}^{-1}$ exists for all $l$ and $\rho\in[0,1/2)$ and $J_{m}(S)$ is an unbiased estimate of $I_{G}(S)$. The next lemma helps us to bound the error of the estimate.
Lemma 6.4
Given $1\leq l\leq k$, for all $C_{\rho,l}$ with $0\leq\rho<1/2$ defined in (3), as $\rho\to 1/2$, $|1-C_{\rho,l}^{-1}(a,b)|=O((1/2-\rho)^{-l^{2}})$ for all $0\leq a,b\leq l$. Thus [^2] Here we first fix $k$ and take $\rho\to 1/2$. Thus, $l\leq k$ are constant relative to $1/2-\rho$., there exists $V_{\rho,k}=O((1/2-\rho)^{-k^{2}})$ where $|1-C_{\rho,l}^{-1}(a,b)|\leq V_{\rho,k}$ for all $l\leq k$ and $a,b\leq l$.
Proof of Lemma 6.4: To simplify the notation, we write $C=C_{\rho,l}\in{\mathbb{R}}^{(l+1)\times(l+1)}$. We will use $a,b$ to denote the indices from 0 to $l$. To bound $|1-C^{-1}(a,b)|$, it is sufficient to bound $|C^{-1}(a,b)|$.
By Cramer’s rule, we can compute the entries of $C^{-1}$ from the ratios of its determinant and its cofactors’ determinants. Specifically, for all $a,b=0,1,\dots,l$, the $(a,b)$ entry of $C^{-1}$ can be obtained as follows: $|C^{-1}(a,b)|=|\det(D^{b,a})/\det(C)|$ where $D^{a,b}$ is called the $(a,b)$ cofactor of $C$, which is a submatrix of $C$ formed by removing the row $a$ and the column $b$ of $C$. Therefore, it is sufficient for us to have an upper bound on $|\det(D^{a,b})|$ for all $a,b$ and a lower bound on $|\det(C)|$.
First, because $C$ is a column-stochastic matrix, $D^{a,b}$ is a column sub-stochastic matrix. We define a diagonal matrix $R\in{\mathbb{R}}^{l\times l}$ where $R_{\ell,\ell}=1/(\sum_{\imath}D^{a,b}_{\imath,\ell})\geq 1$ for all $\ell$ by (3). Because $D^{a,b}R$ is a column-stochastic matrix, $|\det(D^{a,b}R)|\leq 1$ and $|\det(D^{a,b})|\leq|\det(D^{a,b})\det(R)|=|\det(D^{a,b}R)|.$ Therefore, $|\det(D^{a,b})|\leq 1$.
Now we lower-bound $|\det(C_{\rho,l})|$. Because $C(a,b)$ is a degree $l$ polynomial in $\rho$ for all $a,b=0,\dots,l$, $\det(C_{\rho,l})$ is a degree $l^{2}$ polynomial in $\rho$. When $\rho=1/2$, $C_{\rho,l}$ becomes singular, and $\det(C_{1/2,l})=0$, and $\det(C_{\rho,l})\neq 0$ for all $0\leq\rho<1/2$ by LABEL:prop:feasible. Thus, $\det(C_{\rho,l})=\Omega((1/2-\rho)^{l^{2}})$ as $\rho\to 1/2$. The latter, together with the fact that $|\det(D^{a,b})|\leq 1$, implies $|C^{-1}(a,b)|=|\det(D^{b,a})/\det(C)|=O((1/2-\rho)^{-l^{2}})$ which completes the proof. $\Box$
The following is a first-order approximation when $\rho\to 0$ (not used in the paper).
Lemma 6.5
Given $1\leq l\leq k$, for all $C_{\rho,l}$ with $0\leq\rho<1/2$ defined in (3), as $\rho\to 0$, for all $a$
[…]
Therefore, $|1-C_{\rho,l}^{-1}(a,b)|\leq 1+l\rho\leq 1+k\rho$ for all $0\leq a,b\leq l$.
Lemma 6.6
Consider any candidate seed set $S$ of size at most $k$, $\delta>0$, and $m$ perturbed influence samples with $\rho<1/2$. Then $|J_{m}(S)-I_{G}(S)|\leq\delta,$ with probability at least $1-2\exp\left(-{2m\delta^{2}}/{(n^{2}V_{\rho,k}^{2}})\right)$ over the randomness of the perturbed samples, where $V_{\rho,k}=O((1/2-\rho)^{-k^{2}})$ is defined in Lemma 6.4.
Proof of Lemma 6.6: Given $S$ with $|S|=l$, the value of $J_{m}(S)$ is a linear function of the empirical distribution of $\tilde{f}$ on the $m$ samples, so $J_{m}(S)$ can be written as a sum of $m$ iid random variables. Specifically, $J_{m}(S)=\sum_{t=1}^{m}X_{t}$ is a sum of $m$ iid random variable where for $t\in[m]$, $X_{t}=\frac{n}{m}(1-C^{-1}_{\rho,l}(0,a))$ if the intersection of the $t$-th perturbed influence sample and $S$ has size $|\tilde{x}^{t}\cap S|=a$. Thus, the expectation of $J_{m}$ is
$\textstyle\operatorname*{\mathbb{E}}_{\tilde{{\mathbf{x}}}}[J_{m}(S)]$ $=\frac{n}{m}\sum_{t=1}^{m}\sum_{a=0}^{l}(1-C^{-1}_{\rho,l}(0,a))\Pr[|\tilde{x}^{t}\cap S|=a]$
$=$ $n\sum_{a=0}^{l}(1-C^{-1}_{\rho,l}(0,a))\Pr[|\tilde{x}\cap S|=a]$
$=n(1-\Pr[|x\cap S|=0])$ (by definition of $C_{\rho,l}$)
$=n\Pr[x\cap S\neq\emptyset]=I_{G}(S).$
Furthermore, because $|1-C^{-1}_{\rho,l}(0,a)|\leq V_{\rho,k}$, $|X_{i}|\leq\frac{n}{m}V_{\rho,k}$ by Lemma 6.4, we can apply Hoeffding’s inequality hoeffding1994probability to get $\Pr[|J_{m}(S)-I_{G}(S)|\geq\delta]\leq 2\exp\left(-{2m\delta^{2}}/{(n^{2}V_{\rho,k}^{2}})\right),$ for all $\delta>0$, which completes the proof. $\Box$
6.7 Proof of LABEL:thm:rr:acc
Given $\alpha>0$, we set $\delta=\alpha n$ and $m^{*}=(\gamma^{*}{k^{3}V_{\rho_{\epsilon},k}^{2}\ln n})/{\alpha^{2}}$ for some $\gamma^{*}>0$. By LABEL:thm:rr, $I_{G}(\mathcal{M}_{loc}^{k,\epsilon}({\mathbf{x}}))>(1-1/e)\mbox{\sf OPT}-\alpha n,$ with probability
$1-2n^{k}\exp\left(-\frac{m^{*}\alpha^{2}n^{2}}{2k^{2}n^{2}V_{\rho_{\epsilon},k}^{2}}\right)\geq 1-2n^{k}\exp\left(-\frac{\gamma^{*}k\ln n}{2}\right)=1-o(1),$
when the constant $\gamma^{*}$ is large enough.
Run time: Note that the time complexity of computing $\tilde{{\mathbf{f}}}$ in Line 2 of Algorithm 5 is $O(mn)$ and the linear system in Line 3 can be solved using $O(k^{3})$ arithmetic operations, so computing $J_{m}(S_{i-1}\cup\{v\})$ in Line 5 of Algorithm 4 takes $O(k^{3}+nm)$ time. The $nk$ calls to Algorithm 5 by Algorithm 4, bring the overall run time to $O(nk^{4}+kn^{2}m)$. $\Box$
7 Erdős–R’enyi Simulation Experiments and Estimation Details
7.1 Simulations on Synthetic Random Graph Models
To test the performance of our algorithms in a sanitized environment, we generate an input using the Erdős-Rényi random graph model on $200$ nodes with edge probability $0.15$ and run the randomized response algorithm and the exponential mechanism with different privacy budgets ($\epsilon=1,2$, or $3$) and different numbers of influence samples ($m$ ranging from zero to $1000$). The independent cascade probability is set to $0.03$. Figure 7.1 shows that the increase rate with $m$ slows with decreasing privacy budget for both the exponential mechanism and the randomized response algorithm.
Figure 7.1: Performance of our exponential-mechanism and randomized-response algorithms compared with the non-private influence-sample seeding method and the complete-information greedy baseline on an Erdős–Rényi graph with $n=200$ and $p=0.15$. We report the expected influence $I_{G}(\cdot)$ across privacy budgets ($\epsilon$) as the number of influence samples ($m$) increases. Each point represents the mean over $1000$ trials, obtained from $50$ independent collections of $m$ influence samples, with each collection evaluated 2020 times. Error bars show $95\%$ confidence intervals, and the expected spread is estimated using an additional $1000$ influence samples.
As shown in Figure 7.1, both the exponential mechanism and the randomized response algorithm exhibit improved performance as the privacy budget $\epsilon$ and the number of influence samples $m$ increase. When $\epsilon$ is small, the injected noise limits the effective use of network information, resulting in lower expected spread. As $\epsilon\to\infty$, both mechanisms approach the non-private influence-sample greedy baseline, illustrating the trade-off between privacy protection and seeding performance.
7.2 Estimation of STERGM parameters for MSM sexual activity networks
Using ARTnet data, we calculated degree distributions, assortative mixing by age and race, partnership durations by age and race matching, and the probability of concurrent partnerships. These quantities were computed separately for main, casual, and one-time relationships, and we fit a distinct separable temporal ERGM (STERGM) to each. Combining simulated sexual activities from the three STERGMs allowed us to construct MSM sexual activity cascades, which we then used as influence samples for differentially private targeting in Section 4.1. In an STERGM, the transition from a network $Y^{t-1}$ to $Y^{+}$ is modeled through independent formation and dissolution processes, each specified by an exponential random graph model. To fit the STERGMs for the ARTnet data, we use the following formation model:
$\Pr[Y^{+}=y^{+}|Y^{t-1};\bm{\theta}^{+}]\propto\exp\left[\theta^{+}_{e}e+\bm{\theta}^{+}_{d_{a}}\cdot\mathbf{d}_{a}+\bm{\theta}^{+}_{m_{a}}\cdot\mathbf{m}_{a}+{\theta}^{+}_{d_{W}}\cdot{d_{W}}+{\theta}^{+}_{m_{R}}\cdot{m_{R}}+\theta^{+}_{c}c+\bm{\theta}^{+}_{d_{cn}}\cdot\mathbf{d}_{cn}\right],$ (9)
We use bold letters for vectors. Let $e$ denote the total number of partnerships; $\mathbf{d}_{a}$ the mean degree for age group $a$ (15–24, 25–34, 35–44, 45–54, 55–64), with the 15–24 group omitted to avoid multicollinearity when estimating $\bm{\theta}^{+}_{d_{a}}$; and $\mathbf{m}_{a}$ the number of partnerships among individuals in age group $a$. The ARTnet data include two race categories (Black/Hispanic and White/Other), and to prevent multicollinearity we model only one, using $d_{W}$ for the mean degree among White/Other individuals and $m_{R}$ for same-race partnerships. The term $c$ counts individuals with concurrent partnerships. Cross-type degrees $\mathbf{d_{cn}}$ capture the relationship between main and casual partnerships (e.g., individuals with two casual partners have on average 0.36 main partners). To limit cross-type influence, we cap the effect of casual partnerships on main-partnership degree at 3, and the effect of main partnerships on casual-partnership degree at 2. We focus on longer-term relationships and exclude one-off effects on main and casual partnerships (which appear as “–” in Table 1). Based on the formation model in Equation 9, we fit STERGMs to the population-level network statistics from the egocentric ARTnet sample. Dissolution parameters are derived from partnership durations, and, following carnegie2015approximation, we approximate formation parameters by subtracting these dissolution estimates from the ERGM estimates, reducing computational burden for large sparse networks. Parameter estimates for the main, casual, and one-off formation models are shown in Table 1.
Table 1: Parameter Estimates for Main, Casual, and One-off Models
Variable Meaning Main Casual One-off
$\theta^{+}_{e}$ Total number of partnerships -1.4327 (0.5301) -12.3419 (0.7587) -17.0396 (2.5590)
$\theta^{+}_{d_{2}}$ Mean degree of individuals aged 25–34 -1.3258 (0.2866) 0.4676 (0.2536) 3.6471 (0.2119)
$\theta^{+}_{d_{3}}$ Mean degree of individuals aged 35–44 -1.0929 (0.2819) 1.5004 (0.2224) 4.2618 (1.3010)
$\theta^{+}_{d_{4}}$ Mean degree of individuals aged 45–54 -1.3546 (0.2844) 2.2343 (0.2215) 4.5385 (1.2963)
$\theta^{+}_{d_{5}}$ Mean degree of individuals aged 55–64 -1.9009 (0.2918) 2.6267 (0.2294) 4.5570 (1.2855)
$\theta^{+}_{d_{W}}$ Mean degree of White/Other individuals -1.3384 (0.1636) 0.0485 (0.1329) 0.3123 (0.3561)
$\theta^{+}_{m_{1}}$ Partnerships among individuals aged 15–24 0.8243 (0.3111) 3.7552 (0.3465) 9.3462 (2.5612)
$\theta^{+}_{m_{2}}$ Partnerships among individuals aged 25–34 2.2011 (0.3686) 1.4406 (0.2658) 1.8722 (0.5792)
$\theta^{+}_{m_{3}}$ Partnerships among individuals aged 35–44 1.5325 (0.3544) 4.1681 (0.4402) 0.3419 (0.5552)
$\theta^{+}_{m_{4}}$ Partnerships among individuals aged 45–54 1.5553 (0.3903) -0.5821 (0.2834) -0.4781 (0.5723)
$\theta^{+}_{m_{5}}$ Partnerships among individuals aged 55–64 1.7164 (0.4268) -1.7033 (0.3133) -0.9140 (0.6603)
$\theta^{+}_{m_{R}}$ Partnerships among individuals of the same race 0.6973 (0.1482) -0.0871 (0.1517) -0.3068 (0.3939)
$\theta^{+}_{c}$ Number of nodes with concurrent relationships -3.1747 (0.2643) 0.4688 (0.1413) -
$\theta^{+}_{d_{cn1}}$ Average cross-network degree (alternative degree=1) -0.9636 (0.1615) 0.9935 (0.3172) -
$\theta^{+}_{d_{cn2}}$ Average cross-network degree (alternative degree=2) -1.3927 (0.2006) 0.4984 (0.3189) -
$\theta^{+}_{d_{cn3}}$ Average cross-network degree (alternative degree=3) -1.7150 (0.2493) - -
Note: p<<0.1; p<<0.05; p<<0.01
We estimated the generative parameters of sexual behavior for the three partnership types (main, casual, and one-off) using separate STERGMs informed by the previously calculated network measures. These models were then used to generate influence samples that trace direct and indirect sexual relationships over time—reflecting realistic settings where such samples are observable, while the full underlying network is only hypothetical and impractical (and undesirable) to collect. In our simulation of 1,000 individuals over 120 weekly periods, grouped into 10 quarters, we selected 150 initial nodes per quarter and identified all individuals directly or indirectly connected through sexual activity, producing 1,500 influence samples. A second set of 1,500 samples, generated by choosing different initial nodes, served as a test set for evaluating the seed sets produced by our DP seeding algorithms.
7.3 Privacy Semantics and Extensions of ISDP
To clarify the privacy notion in our framework, ISDP protects activity-level participation—whether an individual takes part in a specific diffusion event. Each influence sample represents one observed cascade and records who becomes involved; aggregating such samples yields a dataset of individuals’ participation in observable diffusion activities. ISDP therefore safeguards these participation records rather than structural connections in an underlying social network. In contrast, traditional node- or edge-level differential privacy assumes a fully known network whose nodes or edges can be perturbed, which is unrealistic in practice and would require adding far more noise, severely degrading performance. Although one could in principle adapt our exponential mechanism to satisfy node- or edge-level DP by scaling the noise from $\epsilon$ to $\epsilon/m$, this would substantially reduce utility.
Proposition 7.1 (label=prop:reduction, name=Reduction to Node and Edge DP)
Given $k\leq n\in\mathbb{N}$, $\epsilon>0$, and $m\in\mathbb{N}$ influence samples $\mathbf{x}$}, $\mathcal{M}^{k,\epsilon}_{\mathrm{exp}}({\mathbf{x}})$ in Algorithm 3 is $\frac{\epsilon}{m}$-node DP and $\frac{\epsilon}{m}$-edge DP.
Proof.
Proof of LABEL:prop:reduction Let $\Phi:\mathbb{G}\!\to\!\mathcal{X}$ map a network $\mathcal{G}\in\mathbb{G}$ to its $m$ influence samples ${\mathbf{x}}=\Phi(\mathcal{G})$. Under ISDP, two datasets ${\mathbf{x}},{\mathbf{x}}^{\prime}$ differ in one individual’s participation in one sample, whereas under node- or edge-level DP, $\mathcal{G},\mathcal{G}^{\prime}$ differ by a single node (and its incident edges) or a single edge. For the exponential mechanism with utility $u({\mathbf{x}},v)=\tfrac{n}{m}|C_{{\mathbf{x}}}(S\!\cup\!\{v\})\setminus C_{{\mathbf{x}}}(S)|$ for $v\in[n]\setminus S$, the global sensitivity $\Delta(u)=\max_{Z\sim Z^{\prime}}\max_{v}|u(Z,v)-u(Z^{\prime},v)|$ satisfies $\Delta_{\text{ISDP}}(u)=n/m$ because changing one entry in $\mathbf{x}$} affects at most one sample. In contrast, modifying one node or edge can alter the outcomes of all $m$ samples (e.g., in a star graph), yielding $\Delta_{\text{node/edge}}(u)=n$. Since the exponential mechanism guarantees $\epsilon$-DP with exponent $\exp(\epsilon u/2\Delta(u))$, maintaining the same privacy level under node- or edge-level adjacency requires $\epsilon_{\text{node/edge}}=\epsilon_{\text{ISDP}}\Delta_{\text{ISDP}}(u)/\Delta_{\text{node/edge}}(u)=\epsilon_{\text{ISDP}}/m$. Hence, an $\epsilon$-ISDP mechanism satisfies node- or edge-level DP only with effective budget $\epsilon/m$. $\Box$
Note that to achieve $\epsilon$-node or edge-level DP, we need to run $\mathcal{M}^{k,\epsilon/m}_{exp}$ and the upper bound of utility loss $(1-e^{-1})I_{\mathbf{x}}(S^{*})-I_{\mathbf{x}}(\mathcal{M}^{k,\epsilon/m}_{exp}({\mathbf{x}}))\leq 2k^{2}n(\ln n+t)$ is independent of the number of influence sample $m$ by LABEL:thm:exp:inf. This align with Propositions 2.5 and 2.6 because node- or edge- DP’s protection should be invariant under the number of influence samples. On the other hand, increasing the number of influence sample only reduce the estimation error between $I_{\mathbf{x}}(S^{*})$ and $I_{\mathcal{G}}(S^{*})$ but does not shrink the DP‑induced optimization gap. Finally, while mechanisms other than Algorithm 3 may be used for node‑level or edge‑level privacy, Propositions 2.5 and 2.6 imply that the same utility gap persists.
Another natural extension of ISDP is individual-level ISDP, where individuals may participate in multiple cascades—for example, frequent contacts or high-risk participants. In this setting, changing one individual’s participation affects several samples, corresponding to a full column in the influence-sample matrix. To guarantee privacy across cascades, the privacy budget $\epsilon$ is conservatively scaled to $\epsilon/m$, where $m$ denotes the number of influence samples. This scaling is a worst-case bound, since in practice a person rarely appears in all $m$ cascades; the noise magnitude can therefore be adjusted based on empirical participation frequencies.
Overall, ISDP captures privacy at the level of observable diffusion activities rather than latent network structures, providing a realistic and flexible framework for privacy protection while preserving both analytical tractability and algorithmic utility.
[…]
Liu, Rahimian, and Yu contributed equally and are listed alphabetically. Rahimian is partially supported by the National Science Foundation under Grant No. 2318844 and acknowledges support from a Pitt Momentum Funds Grant on Socially Responsible Data Collection and Network Intervention Designs. This research was supported in part by the University of Pittsburgh Center for Research Computing, RRID:SCR_022735, through the resources provided. Specifically, this work used the H2P cluster, which is supported by the NSF award number OAC-2117681. Data and code for reported simulations can be accessed from https://github.com/aminrahimian/dp-inf-max.